### v2

# A cross-platform analysis of polarization and echo chambers in climate change discussions

Abstract
With the intensification of climate change discussion, social media has become prominent in disseminating reliable and unreliable content. In this study, we present a cross-platform analysis on Youtube and Twitter, and examine the polarization and echo chambers in social media discussions in four datasets related to climate change: COP27, IPCC, Climate Refugees, and Doñana. We have identified communities of users spreading misinformation on Twitter, although they remain relatively isolated from the rest of the network. The analysis by interaction type reveals that climate change sceptics use mentions to draw the attention of other communities. The YouTube posts referenced on Twitter reveal a strong correlation in the community organisation of social media, suggesting a platform alignment. Moreover, we report the presence of echo chambers in YouTube post-sharing related to mainstream and sceptical content.
While a certain level of societal conflict can benefit democracies, an excessively polarized social environment can detrimentally affect their self-governing capacity [1], especially when it pertains to issues that require a high level of consensus to be addressed. The convolution of a highly polarized social environment and an overload of fast-spreading controversial or unreliable information impacts public perception of reality [2]. This dynamic, exacerbated during emergencies and crises [3, 4], further complicates the quest for solutions, rendering resolution increasingly challenging.
A social system is considered to be maximally polarized when the population is divided into two groups of similar size with diametrically opposed views that remain almost isolated [5, 6]. Studies and experiments across various platforms [7, 8, 9] have shown an increase in polarization attributed to the use of social media. However, the precise impact of online social networks on public debate remains ambiguous [10]. Evidence might vary across topics of discussion and online platforms. While Twitter tends to display high polarization, other platforms, such as WhatsApp, have even shown de-polarization [11]. The analysis of Twitter debates across different topics [12] revealed that the discussion on Obamacare was far less polarized than the debate on abortion. Other platforms such as YouTube or Instagram also display polarization [13, 14, 15, 16] Additionally, external events have been observed to heighten activity and polarization on social media platforms [17, 18, 19].
Even though the interest in measuring social and political polarization dates from before the digital era [20, 6], the widespread of online social platforms and data accessibility sparked the development of measures to quantify it. There are two main categories of polarization metrics, along with combinations of them, employed in the analysis of digital traces within online social platforms: those centred on textual analysis [21], and those focusing on interactions and network structure [22].
The first metrics developed to quantify the macroscopic level of structural polarization primarily concentrate on the modularity and the analysis of community boundaries [23, 24, 22]. However, these metrics only provide a partial perspective on polarization, as they largely disregard the multi-scalar organisation of networks and the role of individual nodes. More refined metrics, which capture long-range structural properties, include label propagation [25], centrality measures, and random walk diffusion [26]. Recent studies have aimed to integrate the analysis of structural polarization with individual opinions [12]. Even though most of the measures assume a dipole organisation, there have been recent developments to quantify multipole polarization [27].
Several studies indicate that a phenomenon closely intertwined with the rise of polarization in social media is the prevalence of echo chambers [28, 29, 30], where interactions predominantly occur among individuals who share similar beliefs and ideas. Research exploring the existence and strength of echo chambers across various social media platforms has revealed a spectrum of prevalence and intensity [29, 7, 31, 14, 32, 33]. Notably, platforms like Facebook and Twitter exhibit more pronounced echo chambers compared to Gab and Reddit. The presence of echo chambers can be quantified through the political leaning of an individual and its neighbourhood [34]. As interest in modelling the dynamics of polarization and echo chambers grows [35, 36], efforts to discern the multiple factors influencing social media interaction patterns face significant challenges. Even the recommendation algorithms employed by these platforms can impact observed dynamics [37, 38, 39]. For instance, platforms may inadvertently exacerbate polarization by limiting exposure to opposing viewpoints in users’ feeds [8].
The combination of a polarized environment [40] and psychological traits [41], such as confirmation bias [42, 43, 44], has been closely associated with the dissemination of misinformation [45]. This spread of unreliable and fake news significantly impacts society, influencing outcomes such as presidential elections [46, 47] and vaccine hesitancy [48, 49]. Despite these challenges, there have been concerted efforts to combat the spread of fake news [50]. Nonetheless, debates persist regarding the distinction between reliable and unreliable news cascades. Although initial studies highlighted structural differences [51], subsequent analyses have suggested that disparities may be more attributable to differences in cascade size [52].
Besides politics, climate change stands out among the topics in which unreliable news has proliferated [53, 54]. The existence of climate change has been proven repeatedly from a scientific perspective [55, 56, 57, 58] as has its anthropogenic contribution [59, 60, 61]. There are multiple pieces of evidence on the impact of climate change on health [62], the economy [63], the agriculture [64] or the overall ecological system [65], among others. However, there has been an increasing polarization in social media around the impact of climate change and its anthropogenic nature [66]. Previous research has focused on the origins of this polarization [67], its relation to political affiliations [68], and the influential roles played by media outlets [69] and corporations [70]. Recently, an increasing polarization was reported in locations suffering from extreme weather events [71].
In this work, we analyse the climate change discussion in four datasets from Twitter and YouTube across different contexts. The degree of polarization varies across datasets due to their distinct characteristics and is higher in the Doñana dataset, possibly due to its close relation with local politics. We analysed the interaction between the two largest communities spreading reliable and unreliable content and confirmed a weak interaction between them. Additionally, we provide a cross-platform perspective by examining YouTube links shared on Twitter (currently X) and analysing the bipartite graph derived from these links. Our findings reveal a connection between communities across both platforms, indicating a strong interconnection across social media platforms.
I Data Description
We extracted four datasets related to climate change topics from Twitter through the Twitter Academic API. The first one is related to the 2022 United Nations Climate Change Conference (COP27), the second to the release of the 6th Assessment Report by the Intergovernmental Panel on Climate Change (IPCC), the third to the climate refugee crisis resulting from extreme weather conditions (Climate Refugees), and the fourth to a new legislation concerning the natural reserve of Doñana in the south of Spain (Doñana). The tweet recollection was conducted with Twarc [72] by searching for keywords related to each discussion (Appendix A).
Network Minimum date Maximum date Number of users Tweets Retweets Replies Quotes
COP27 2022-09-01 2022-11-27 1351903 866753 4977874 205973 175078
IPCC 2023-03-18 2023-03-26 157056 31138 267971 45863 7751
Clim. Refugees 2008-03-10 2022-12-31 841454 384267 1376057 139699 38909
Doñana 2019-01-01 2023-04-30 290782 139478 1187646 135245 25056
Table 1: Summary table of the Twitter datasets analysed. Minimum and maximum dates, observed number of users, and number of tweets by typology.
In Table 1 we present the main statistics of the datasets and in Fig. S1 the daily tweet activity, where the Climate Refugees and Doñana datasets exhibit longer timespans than the IPCC and COP27 datasets. The IPCC dataset covers only nine days, coinciding with the heightened discussion surrounding the release of the 6th Assessment Report. We provide in Appendix B the most central users based on the total inflow of retweets (Tables S1, S2, S3, and S4) and the most frequent languages in the ten largest communities of each dataset (Tables S5, S6, S7, and S8).
We extracted the corresponding YouTube dataset by analysing the URLs referenced in each Twitter dataset and downloading the posts and comments using the YouTube Data API. To ensure relevance, we filtered the posts by searching for occurrences of certain keywords in each video title or description (Appendix A).
Network Videos Min. Date Max. Date Avg. views Avg. likes Avg. comments
COP27 624 2008-02-11 2023-05-12 571964.35 5860.32 306.99
IPCC 145 2007-11-03 2023-06-09 350178.66 8991.67 1455.16
Clim. Refugees 215 2009-11-22 2023-11-06 243720.27 3604.81 527.40
Doñana 191 2009-01-11 2023-06-01 101098.32 3622.68 246.25
Table 2: Summary table of the Youtube datasets analysed. Number of videos in each community, minimum and maximum dates observed, and average number of views, likes, and comments.
We present a summary of the YouTube datasets in Table 2, which includes the number of videos, the minimum and maximum dates, and their average statistics. Unlike the Twitter dataset, the period of the comments is longer due to users frequently referencing older videos. Fig. S2 illustrates the time series of comments in each dataset. Each dataset exhibits peaks at different times based on its unique characteristics. In particular, we observe a strong alignment in the IPCC and COP27 datasets with the event dates.
II Twitter analysis
We evaluated the global polarization on Twitter by considering the unweighted network of retweets where links are undirected and indicate a retweet interaction between two users. We used the Metis partition algorithm [73] to split the graph into two clusters, and we calculated the modularity [74, 75], the E-I index [76] and the adaptive E-I index [77] which accounts for the uneven size of clusters. The results for the four datasets analysed are presented in Fig. 1. Following the methodology proposed in [22], we computed polarization metrics in the observed network $\Phi(G)$, as well as in the ensemble of configuration models with equivalent degree distribution $\Phi(G_{CM})$ [78], and calculated the denoised value $\hat{\Phi}(G)=\Phi(G)-\Phi(G_{CM})$.
Figure 1: Polarization in the Twitter dataset. Quantification of the polarization in Twitter: (a) modularity, (b) E-I index, and (c) adaptive E-I index. We provide the measures for the observed networks in red, across 100 realizations of the configuration model with preserved degree sequences in blue, and the difference between them (denoised value) in yellow.
All networks exhibit polarization according to the denoised value. In particular, the COP27 and Doñana datasets displayed larger denoised values, suggesting a stronger polarization around the topics. The higher polarization observed in the COP27 dataset may be attributed to the presence of communities using different languages that do not interact with each other (Table S5). The Doñana dataset captures the discussion around a new law related to natural park preservation, which involves political parties and users with strong political affiliations. The lower polarization observed in the IPCC dataset may be due to its nature as a scientific report release rather than an international conference, potentially attracting a more scientifically educated audience and inducing a lower level of polarization. The standardised values [22] (Fig. S3) indicate that polarization is stronger in the Climate Refugees dataset.
We examine the network of retweets and mentions between the $10$ largest communities of the networks in Fig. 2. We identified these communities using modularity optimisation on the directed weighted graph of retweets [75]. The interaction patterns between the largest community (labelled as 0) and the remaining communities vary across datasets and types of interactions. In the IPCC and COP27 networks, the smaller communities mostly retweet content from the largest community, which is more isolated and engages in fewer interactions with the rest of the network. Those largest communities feature official accounts such as the Secretary General of the United Nations or the Intergovernmental Panel on Climate Change (Tables S1 and S2). Conversely, the Doñana network exhibits greater flow heterogeneity between communities, with community 5 playing a central role in attracting most of the flow. The detailed overview of the users in the community reveals the presence of the president of Spain and the ruling political party (Table S4). There are significant structural differences depending on the type of interaction. For instance, in the IPCC network, there are relatively few retweets from community 1 to community 0, but there is a high volume of mentions.
Figure 2: Community interaction network. Interaction between the ten largest communities in (a) COP27, (b) IPCC, (c) Climate Refugees, (d) Doñana. Arrows indicate the direction of the interaction, the colour and width, and the volume. The size of the dots represents community size. In red, we show the interaction through retweets and, in blue, through mentions.
To assess the political bias and reliability of each community, we calculated the user values by averaging over the URLS they shared. Information on the political bias and reliability of news outlets was obtained from MediaBias fact-check [79], where positive and negative political bias correspond to right-wing and left-wing media, respectively. Outlets with higher reliability are considered more trustworthy and tend to share more verified news. For the Doñana dataset, the reliability and bias were based on a separate dataset [80], given that most media sources are Spanish. To facilitate the analysis, we have normalised the bias within the range [-1, 1] and the reliability within the range [0, 1]. The minimum and maximum bias are set according to the largest absolute value observed. In Fig. 3(a)-(d), we show the density plot of community reliability as a function of their political bias according to the shared URLs in the IPCC, COP27, Climate Refugees, and Doñana networks. We computed user-level values by averaging over the sources they shared, and then calculated community-level values by averaging over the users. All datasets have a left bias as most communities display negative values, particularly pronounced in the IPCC and Climate Refugee networks. Possible reasons for this include bias in the source data, climate change sceptics sharing non-official sources, or an over-representation of left-biased users in climate change discussions. Left-wing users might interact more in the context of climate change due to their higher awareness. The left bias is particularly pronounced in the IPCC and Climate Refugee networks, where the cluster of large communities is centred around negative values. However, the bias is more dispersed in the Doñana dataset, possibly due to differences in the shared media sources or a wider ideological variety. We show the user distributions of bias and reliability for the seven largest communities in Fig. 3(e)-(h) with most of them peaked at negative values (left bias). The main communities with right bias we have identified are community one in the COP27 and IPCC, community six in Climate Refugees, and community zero in Doñana.
Figure 3: Reliability as a function of the political bias in the networks studied. (a-d) Density plot of the distribution of reliability as a function of the bias for the communities detected in (a) COP27, (b) IPCC, (c) Climate Refugees and (d) Doñana. (e-h) Distribution of bias and reliability in the seven largest communities for (e) COP27, (f) IPCC, (g) Climate Refugees and (h) Doñana.
Based on the previous findings and manual inspection of the communities (See Tables S1, S2, S3, and S4), we identified the two largest communities with opposing political leanings. In the IPCC and COP27 datasets, community 0 represents the largest community aligned with the mainstream climate change narrative, while community 1 is the primary advocate of a counter-narrative message. However, in the Doñana network, the labels are swapped because the right-biased community has a larger user base. For the Climate Refugees dataset, most of the top 10 communities exhibit similar bias-reliability values, except for community 6, which stands out with low reliability and a right bias. In this dataset, we designate community 1 as the primary left-biased community, as it includes well-known influencers such as Greta Thunberg and engages in more interactions with the sceptic community.
To evaluate the isolation of the communities, we calculated the directed interactions between them, and we computed the normalised outflow difference given by
$P^{out}_{i}=\frac{\sum_{j,j\neq i}T_{ij}-T_{ii}}{\sum_{\forall j}T_{ij}},$ (1)
and the entropy of flows as
$H^{out}_{i}=-\sum_{\forall j}p_{ij}logp_{ij},$ (2)
where $T_{ij}$ is the number of interactions from community $i$ to community $j$ and $p_{ij}$ is the probability of having a retweet from $i$ to $j$ given by $T_{ij}/\sum_{\forall j}T_{ij}$. While the first index quantifies the ratio between external and internal links and goes from $-1$ to $1$ when all flows are external, the entropy provides information on the variety of those flows. The entropy is 00 when there is no variety of flows and $1$ when it is maximum.
In Fig. 4, we represent the outflow entropy as a function of the isolation for the four networks. Most of the largest communities exhibit a certain degree of isolation, indicated by negative values. Across all datasets, the right-biased community demonstrates strong isolation and low entropy, implying their messages have limited spread across the network. However, the pattern changes when considering the mention graph (Fig. S4). The largest mainstream communities show minimal mentions of other communities, whereas the low-reliability communities display lower isolation and more external mentions. This observation suggests that climate change sceptics may reference the mainstream community to attract attention but get little attention instead.
Figure 4: Community connectivity and entropy in retweet networks. Outflow entropy as a function of the normalised outgoing isolation in the retweet networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of feasible values.
We computed the minimum and maximum attainable entropy values as a function of $P^{out}_{i}$ and $p_{ii}$ for orientation. Given a value of $P^{out}_{i}$ we have
$\frac{\sum_{j,j\neq i}T_{ij}}{{\sum_{\forall j}}T_{ji}}=P_{i}+\frac{T_{ii}}{{\sum_{\forall j}T_{ij}}}=P_{i}+p_{ii}.$ (3)
The minimum value of the entropy corresponds to the situation when a community is connected only to itself and one external community (except for the extreme case $P_{i}=1$), which we can plug into the entropy as
$H^{in}_{i}=-p_{ii}logp_{ii}-P_{i}+p_{ii}logP_{i}+p_{ii}.$ (4)
The maximum value of the entropy appears when a community is linked to all the others with a flow $T_{ij}/(N_{c}-1)$, where $N_{c}$ is the total number of communities in the network. Thus, it can be written as
$H^{in}_{i}=-p_{ii}logp_{ii}-\sum^{N_{c}-1}_{j=0}\frac{P_{i}+p_{ii}}{N_{c}-1}log\frac{P_{i}+p_{ii}}{N_{c}-1}.$ (5)
In Fig. S5, we present the reversed quantities computed for the inflow, where we also observe that the right-biased communities with low reliability are among the most isolated.
Given that many networks include users from different countries speaking various languages, we also calculated polarization by focusing on the two largest communities speaking the same language with opposing biases, comparing the networks of retweets and mentions (Fig. S6). They still exhibit polarization in the retweet network, however, the scores are considerably smaller or even negative in the mention network. Thus, we confirm the previous findings on the disparity in network organization based on interaction typology. Our results suggest that focusing solely on retweet networks may provide a biased perspective on Twitter dynamics and that effective polarization, in terms of interaction between opposing views, may be low.
II.1 User analysis
Throughout this section, we assess the presence of echo chambers with a focus on the two largest communities exhibiting opposing political biases. We employed an approach developed in [31] to evaluate the existence of echo chambers. For a given user $i$ with degree $k_{i}$, we computed its average bias $x_{i}$ and reliability $y_{i}$, along with the average bias and reliability of its neighbourhood, denoted as $\frac{1}{k_{i}}\sum_{\forall j\neq i}A_{ij}x_{j}$ and $\frac{1}{k_{i}}\sum_{\forall j\neq i}A_{ij}y_{j}$, respectively.
Figure 5: Detection of bias echo chambers in the retweet networks. Neighbourhood political bias as a function of the user bias in the retweet network for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
In Fig. 5, we show the density plot for the bias of the neighbourhood of users as a function of their own values when considering the interactions through retweets. As expected from the overall ideological positioning of the users, there is a hotspot on the left of the ideological spectrum and a weaker one on the right side. We have identified a diagonal trend that could be related to the presence of echo chambers. Right-biased users interact more with left-biased users, as there are two hotspots for positive values in some cases. The presence of echo chambers is more evident in the Doñana dataset than in the rest of the networks, mainly due to the presence of more users with right bias. There are three hotspots of comparable magnitude across the diagonal with right, centre, and left bias. Still, the hotspot on the left side of the spectrum displays a lower ideological isolation, since it interacts with users featuring a bias close to 0. The right-biased users interact mainly between themselves and not with users in the centre. When analysing other types of interactions, such as the mentions (Fig. S7), the presence of echo chambers is weaker with nearly horizontal hotspots in the COP27 and IPCC datasets. This observation is consistent with the results on polarization from the previous section. The existence of echo chambers is less pronounced when considering different types of interactions, and mentions could serve as a bridge between users with different political biases.
We also examine the echo chambers of the content reliability in Fig. S8, where there are three hotspots for high, medium, and small values. The reliability has a similar behaviour, the hotspots in high reliability suggest that users who share highly reliable sources interact with each other, while low-reliability users interact with users that spread highly reliable sources.
Figure 6: Chamber overlap distribution $q_{ij}$. Chamber overlap distribution for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana calculated on the top $20$ users of each community.
We inspect in detail the user interaction similarity by quantifying their echo chamber [81], without considering the ideology. Given a set of leading users $i$, which have an audience $A_{i}$ composed of the set of users that retweeted it, their chamber $C_{i}$ is the set of users retweeted by audience $A_{i}$. We focus on the two communities with opposing political biases, specifically the top 20 users with the most retweets in each of those communities. The first quantity we will analyse is the chamber overlap between a user $i$ and a user $j$ given by
$q_{ij}=\frac{C_{i}\cap C_{j}}{C_{i}\cup C_{j}},$ (6)
where $C_{i}$ and $C_{j}$ are the chambers of users $i$ and $j$. It is worth noting that the quantity $q_{ij}$ is symmetric by definition and is equal to $1$ when chambers are identical and 00 when they are completely different.
In Fig. 6, we show the chamber overlap distribution $P(q_{ij})$, where a notable disparity between networks is observed. A strong bimodal distribution suggests the presence of echo chambers, as each user has low values of $q_{ij}$ with users outside its echo chamber and high values with those inside it. The results from the COP27 and Climate Refugees datasets suggest that leaders have more heterogeneous chambers compared to the IPCC and Doñana datasets, where we observe a peak for high chamber overlap. The Doñana dataset features a strong echo chamber effect with two separated peaks at low and high values of $q_{ij}$. In Fig. S9, we have performed hierarchical clustering of the top 20 users per community based on the similarity of their chambers, showing a strong alignment between the communities detected.
II.2 Hashtag analysis
Next, we inspect the patterns of hashtags shared in each network, focusing on the two largest communities with opposing biases. We report in Fig. S10 the top hashtags per network, where we show that despite the ideological confrontation, they share a large number of them. The mainstream communities use rather generalistic hashtags such as #$\#COP27$, #$\#IPCC$, #$\#climatechange$, #$\#ar6$, or #$\#climatereport$. The results suggest that the users of the right-biased community use common hashtags to attract the attention of mainstream users. In the case of the Doñana, we observe hashtags related to the environment as #$\#donana$, #$\#donanaseextingue$, or #$\#salvemosdonana$ for the mainstream community. The community with right bias has hashtags related to right-wing parties as #$\#teamvox$, or against the left-wing government as #$\#gobiernodimision$. Common hashtags between communities are less frequent in the Doñana dataset in line with the more pronounced echo chambers detected.
To compute the hashtag polarization, we calculated the fraction of tweets with each hashtag by community given by
$P_{i}^{\#}=\frac{N_{i}^{r}-N_{i}^{l}}{N^{l}_{i}+N_{i}^{r}},$ (7)
where $N_{i}^{l}$ and $N_{i}^{r}$ are the number of tweets with the hashtag $i$ in the main left and right-wing communities, respectively. We display the polarization of the top hashtags in Fig. 7. The number of hashtags in each network varies since we have considered the $20$ most shared in each community, and some of them can overlap. Interestingly, we observe only a few hashtags with values close to 00, suggesting a strong polarization in the use of hashtags. In the IPCC and COP27 datasets, the mainstream community dominates most of the hashtags, while the community spreading a counter-narrative dominates only a few hashtags related to the denial of climate change that include #$\#climatehoax$ or #$\#agenda2030$. In COP27, we also observe hashtags related to hypocrisy as #$\#greenwash$. In the Doñana network, we also observe only a few hashtags close to zero and hashtags almost equally divided between $+1$ and $-1$. We have analysed the chamber overlap of hashtag dynamics in Fig. S11.
Figure 7: Hashtag polarization. Polarization around hashtags for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. Negative and positive values correspond to hashtags more abundant in the left-biased and right-biased communities, respectively.
III Youtube analysis
We have analysed the dataset created from the YouTube videos quoted and referenced on Twitter. The information gathered for each post includes the transcription and description of each video, the corresponding channel, and the user comments. The comment information allows us to create a bipartite network between users and posts since we know the users who have commented on each video. We can project the bipartite network into a post network or a user network. In the post network, the weight corresponds to the number of users in common and, in the user network, to the posts in common. The graphs are undirected by construction in both cases. In Fig. S12, we display the network between YouTube posts for each dataset. We performed a community analysis using the greedy optimisation of modularity on the weighted network [75]. The organisation of communities is complex and fuzzy with several interconnections except for the Doñana network, where communities 00 and $1$ have a clear separation. By inspecting the videos with a large weighted degree, we have identified the communities with mainly climate change sceptics: community $2$ in COP27, communities $1$ and $2$ in IPCC, community $2$ in Climate Refugees, and community 00 in Doñana (Fig. S13).
Despite the communities being more separated, there is a strong interaction between them, indicating a certain level of engagement with opposed content. The exception is the Doñana network, where communities exhibit a clear separation, aligning with the pronounced polarization observed in the corresponding Twitter network. We computed polarization indices on the subnetworks using modularity and EI indices, similar to our approach for Twitter (Fig. 8). Only the COP27 and Doñana networks show higher levels of polarization compared to the others. Overall, despite confronting views on climate change, the interaction between posts remains robust. We conducted a sentiment and emotion analysis of the comments by the community (Fig. 9), revealing that negative comments dominate all posts. However, the posts related to sceptic content feature higher negative sentiments, together with more anger and disgust emotions. The sentiment and emotions of comments were calculated following [82]
Figure 8: Polarization in the Youtube dataset. Quantification of polarization in the post network through (a) modularity (b) E-I index and (c) adaptive E-I index. We have calculated the indices in the real networks and in 100 randomisations where we preserve the degree sequences.
Figure 9: Average sentiment of comments in YouTube posts by community. Average sentiment and emotions: (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana split by mainstream and sceptic communities.
We projected the bipartite graph into a user network to analyse the system from an individual interaction perspective. In Fig. S14, we plot the user interaction, where we focus on the comments of users within the posts of the sceptics and mainstream communities that we have previously separated. Each node corresponds to a user, and the colour to the community according to the greedy modularity optimisation. Despite the users being split into several communities, most of them fall into two to three large communities, which suggests a certain level of polarization. For instance, most users fall within the three largest communities for COP27.
We computed the polarization in the user network by splitting it into two parts using the METIS algorithm and calculating the values in the configuration graphs. Unlike in the post networks, we observe clear signs of polarization in all the networks (Fig. S15). Despite the strong interaction between posts with different views, it is clear that the comment patterns of users are highly polarized. If we aim to measure the polarization between individuals, the user network is the best approximation.
Figure 10: Distribution of user ideology in Youtube. Average ideology of the users split between mainstream and sceptic ideas in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. We assigned an ideology to each user according to the most common community of its comments.
We computed the ideology distribution of users by analysing the communities where they post comments. Focusing on the users with four or more comments, we calculated each user’s ideology as the average ideology of the videos they commented on, weighted by the size of each community to address any imbalances. We considered the mainstream community equal to $1$, and the sceptic community equal to 00. Both distributions are normalised separately. Overall, they point out that there is a clear separation in the ideology of users according to the posts where they have commented.
IV Cross-platform analysis
Finally, we conducted an analysis linking Twitter and YouTube social media platforms. We computed the frequency with which each Twitter community references YouTube posts and their corresponding communities. To address the heterogeneity in community sizes observed in the YouTube dataset, we normalised the number of references to each YouTube community by the community size. Fig. 11 illustrates a strong connection between Twitter and YouTube communities, as most references in each Twitter community align with a specific YouTube community. Additionally, each Twitter community demonstrates a unique pattern of references to YouTube.
In the COP27 network, the larger left-biased Twitter community predominantly references the YouTube community 00, while the larger right-biased community solely references posts in the YouTube community $1$. Conversely, in the IPCC network, the larger right-biased community references various YouTube communities, indicating that sceptic YouTube posts are dispersed across different communities. Except for the IPCC network, posts shared by the larger right-biased community on Twitter predominantly align with a single YouTube community. Furthermore, sceptic posts on YouTube receive minimal references from other Twitter communities.
Figure 11: Interaction between Twitter and Youtube communities. Number of references to YouTube communities by the top 10 Twitter communities in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. To account for the uneven size of YouTube communities, we have divided the counts by the YouTube community size.
To inspect in detail the patterns of YouTube post sharing, we conducted a similar analysis to the echo chamber methodology developed for the Twitter datasets. Using the community assignment of posts between sceptics and mainstream, labelled as 00 and $1$ respectively, we computed the average ideology of users and their neighbourhoods. The density plot in Fig. 12 has clearly defined echo chambers, with two centres on opposite sides of the spectrum. These results suggest that users close to each other share similar content. However, there are differences across the typology of videos shared, as users who share sceptical videos also share mainstream videos to some extent.
Figure 12: Twitter echo chambers in the sharing of YouTube links. Density plots of the average ideology of the videos shared by the neighbours of a user as a function of its average ideology.
V Discussion
In this study, we have assessed the dynamics of polarization and echo chambers on climate change discussions around four different events. Our findings reveal that the extent and nature of polarization can vary significantly depending not only on the specific topic but also on contextual factors. Interestingly, we noted an over-representation of left-biased media in all datasets except one, indicating the dominance of left-leaning perspectives in climate change discussions, possibly affecting media attention from right-aligned outlets. In discussions more directly related to partisan political contexts, such as Doñana and COP27, we observed a stronger polarization and distinct echo chambers than the discussion focused on factual information, centred around the IPCC’s Report. This suggests that proximity to partisan politics amplifies polarization and the formation of echo chambers within online climate change discourse. This may indicate a dynamic of partisan sorting, as already found in the Finnish Twittersphere [77]. Our analysis of echo chambers revealed nuanced findings, particularly regarding interaction types. While the networks of user mentions in COP27 and IPCC discussions showed minimal evidence of echo chambers, retweet networks prominently featured echo chamber dynamics, indicating a higher level of endorsement associated with retweets compared to other interaction types. Expanding our analysis to include YouTube videos referenced in Twitter datasets provided a cross-platform perspective on climate change discourse. We identified unique communities within YouTube videos, separating sceptic perspectives from mainstream content. Interestingly, alignment between Twitter and YouTube communities is strong, with Twitter communities exhibiting distinct preferences for sharing YouTube content, especially for the sceptic community. Thus, indicating the presence of cross-platform echo chambers. Our work suggests that polarization is ubiquitous across social media platforms, although the typology of interactions impacts the overall polarization.
Data availability statement
The Twitter data are made available in accordance with Twitter’s terms of service at https://github.com/clint-project/clint-data and the codes to calculate the metrics at https://github.com/clint-project/polarization_metrics.
Author contribution
A.B: data collection, data curation, investigation, methodology, visualization; J.M.: data collection, data curation, investigation, methodology. E.C.: conceptualization, investigation, methodology, writing; J.V.: conceptualization, investigation, methodology, writing. All authors gave final approval for publication and agreed to be held accountable for the work performed therein.
Acknowledgements
J.M. was a fellow of Eurecat’s “Vicente López” PhD grant program. This work was financially supported by the Catalan Government through the funding grant ACCIÓ-Eurecat (Project TRAÇA 2023 - CLINT). E.C. acknowledges support from the Spanish grants PGC2018-094754-B-C22 and PID2021-128005NB-C22, funded by MCIN/AEL 10.13039/501100011033 and ”ERDF A way of making Europe”, and from Generalitat de Catalunya under project 2021-SGR-00856.
Appendix A Dataset description
In this paper, we employed Twarc, a tool for collecting data from Twitter, to delve into user-generated content. By utilising hashtags as search queries, particularly the search and stream API, we collected Twitter conversations, targeting specific themes and topics relevant to our study (COP27, IPCC, Climate Refugees, and Doñana). This methodological approach facilitated the systematic collection of rich and diverse data. These are the specific hashtags that we utilised to build the datasets:
  • COP27: COP27, TogetherForImplementation, COPTV, PRECOP, PRECOP27, LossAndDamage, Egypt_COP27, COP27Egypt, ClimateAction, ClimateCrisis, ClimateJustice.
  • IPCC: ipcc, @ipcc_ch.
  • Climate Refugees: climate (refugees OR refugee), climate migration, climate displaced.
  • Doñana: Doñana.
  • COP27: COP27, climate, COP
  • IPCC: IPCC, report, climate.
  • Climate Refugees: Refugee, climate.
  • Doñana: Doñana, clima.
In Fig. S1 and Fig. S2 we show the timeseries of each dataset for Twitter and YouTube, respectively.
Figure S1: Timeseries of the COP27, IPCC, Climate Refugees and Doñana datasets. Daily timeseries of tweets in the datasets gathered for COP27, IPCC, Climate Refugees and Doñana.
Figure S2: Timeseries of comments in the Youtube datasets. Number of comments per date in the Youtube datasets of COP27, IPCC, Climate Refugees and Doñana. For visualization purposes, we show the comments after 2021.
Appendix B Influencers by community
We report in Tables S1, S2, S3 and S4 the top influencers by community for the COP27, IPCC, Climate Refugees, and Doñana datasets, respectively.
User Community Pagerank Inflow In-degree Outflow Out-degree Flow ratio Degree ratio
MikeHudema 0 0.004604 126323 42099 12 8 0.999905 0.999810
COP27P 0 0.008392 108582 45664 380 142 0.996513 0.996900
antonioguterres 0 0.006937 79507 38380 5 5 0.999937 0.999870
UN 0 0.003134 43045 24757 211 56 0.995122 0.997743
JamesMelville 1 0.007062 49172 31076 8 5 0.999837 0.999839
DiEM_25 1 0.005087 28848 28496 11 6 0.999619 0.999789
BernieSpofforth 1 0.002298 26501 15218 21 10 0.999208 0.999343
DrEliDavid 1 0.000952 13377 11137 0 0 1.000000 1.000000
kimpaim 2 0.001172 24266 18407 2 2 0.999918 0.999891
TerraBrasilnot 2 0.001122 16518 12962 0 0 1.000000 1.000000
VemPraRua_br 2 0.000899 16234 13991 0 0 1.000000 1.000000
marciolabre 2 0.000652 13147 13028 0 0 1.000000 1.000000
LulaOficial 3 0.010655 130064 38689 24 10 0.999816 0.999742
choquei 3 0.005587 25475 14922 1 1 0.999961 0.999933
Reuters 3 0.001247 12269 8240 67 13 0.994569 0.998425
MarinaSilva 3 0.002625 11904 8575 11 6 0.999077 0.999301
CarolineLucas 4 0.001137 33989 14931 40 31 0.998825 0.997928
NicolaSturgeon 4 0.000565 10778 4993 59 25 0.994556 0.995018
_david_ho_ 4 0.000498 6706 6246 8 8 0.998808 0.998721
johnestevens 4 0.000809 6553 6154 2 2 0.999695 0.999675
RishiSunak 5 0.005127 37770 24640 3 3 0.999921 0.999878
EmmanuelMacron 5 0.003232 16585 12178 2 2 0.999879 0.999836
Poulin2012 5 0.001389 12919 12735 3 3 0.999768 0.999764
MickaCorreia 5 0.000466 3869 3054 45 19 0.988503 0.993817
POTUS 6 0.004068 24808 15824 1 1 0.999960 0.999937
SpeakerPelosi 6 0.000971 9494 5361 9 6 0.999053 0.998882
JustinTrudeau 6 0.000316 5750 4247 2 2 0.999652 0.999529
washingtonpost 6 0.001279 5564 4655 14 2 0.997490 0.999571
Novozymes 7 0.000194 7735 3629 16 10 0.997936 0.997252
AfDB_Group 7 0.000376 5991 2294 358 104 0.943613 0.956631
RockefellerFdn 7 0.000438 5845 3381 346 118 0.944112 0.966276
WilliamsRuto 7 0.000691 4941 3163 9 4 0.998182 0.998737
Danielbricen 8 0.001689 12626 7694 4 2 0.999683 0.999740
IvanDuque 8 0.000791 9105 6317 1 1 0.999890 0.999842
alertaLatam 8 0.000337 6695 4717 17 3 0.997467 0.999364
marcelamvzyya 8 0.000140 5918 3507 0 0 1.000000 1.000000
petrogustavo 9 0.005749 54899 24016 23 14 0.999581 0.999417
infopresidencia 9 0.001116 18209 6575 60 17 0.996716 0.997421
FranciaMarquezM 9 0.000399 10795 5306 22 17 0.997966 0.996806
GustavoBolivar 9 0.001154 8534 6545 7 2 0.999180 0.999695
Table S1: Top 5 most relevant users by total inflow in the ten largest communities of the COP27 dataset
User Community Pagerank Inflow In-degree Outflow Out-degree Flow ratio Degree ratio
IPCC_CH 0 0.173282 86702 48491 22 12 0.999746 0.999753
ed_hawkins 0 0.002570 3194 2119 60 29 0.981561 0.986499
MrMatthewTodd 0 0.001954 2766 2474 54 37 0.980851 0.985265
CharlieJGardner 0 0.002120 2720 2297 5 4 0.998165 0.998262
uksciencechief 1 0.004402 12700 5452 1 1 0.999921 0.999817
BernieSpofforth 1 0.005600 5735 4972 0 0 1.000000 1.000000
fmeeus1 1 0.004689 3973 1800 15 8 0.996239 0.995575
Lauratobin1 1 0.000858 3067 1287 4 3 0.998697 0.997674
jasonhickel 2 0.008178 4249 4120 2 2 0.999530 0.999515
greenpeace_esp 2 0.003291 2092 1986 28 11 0.986792 0.994492
millares 2 0.001390 1788 1786 5 4 0.997211 0.997765
ionebelarra 2 0.000286 592 444 3 3 0.994958 0.993289
GretaThunberg 3 0.007815 8539 8115 0 0 1.000000 1.000000
CarolineLucas 3 0.003437 5833 3971 5 5 0.999144 0.998742
paulpowlesland 3 0.001666 1189 1179 0 0 1.000000 1.000000
guardian 3 0.000646 1016 915 0 0 1.000000 1.000000
KHayhoe 4 0.006960 9241 7626 82 39 0.991205 0.994912
SenSchumer 4 0.001571 4078 1973 2 2 0.999510 0.998987
curious_founder 4 0.004436 2067 1779 9 5 0.995665 0.997197
remblance_erin 4 0.001244 1380 1325 9 9 0.993521 0.993253
rahmstorf 5 0.003506 3135 2032 35 17 0.988959 0.991703
micha_bloss 5 0.007133 2687 2276 3 2 0.998885 0.999122
Martin_Bethke 5 0.001183 2124 1222 71 40 0.967654 0.968304
HolzheuStefan 5 0.001040 1787 1241 26 13 0.985659 0.989633
antonioguterres 6 0.015073 6512 4285 1 1 0.999846 0.999767
UN 6 0.007672 4093 2753 8 2 0.998049 0.999274
BarackObama 6 0.002001 1794 1775 0 0 1.000000 1.000000
HillaryClinton 6 0.001469 1398 1375 0 0 1.000000 1.000000
JKSteinberger 7 0.006622 2121 1930 76 26 0.965407 0.986708
frantecol 7 0.001867 2101 2052 6 4 0.997152 0.998054
jennystojkovic 7 0.000888 1519 1324 8 4 0.994761 0.996988
NLRebellion 7 0.000375 571 407 26 17 0.956449 0.959906
MikeHudema 8 0.003279 5639 3757 1 1 0.999823 0.999734
s_guilbeault 8 0.001042 2369 1082 4 2 0.998314 0.998155
tveitdal 8 0.000368 561 409 35 7 0.941275 0.983173
gmbutts 8 0.000383 512 406 4 3 0.992248 0.992665
DrLuetke 9 0.000737 741 719 0 0 1.000000 1.000000
Axel_Bojanowski 9 0.001738 561 406 11 3 0.980769 0.992665
welt 9 0.000674 543 471 2 1 0.996330 0.997881
Schlautropf 9 0.000378 419 375 14 12 0.967667 0.968992
Table S2: Top 5 most relevant users by total inflow in the ten largest communities of the IPCC dataset
User Community Pagerank Inflow In-degree Outflow Out-degree Flow ratio Degree ratio
AOC 0 0.012350 35741 31238 9 7 0.999748 0.999776
BernieSanders 0 0.002378 12756 11197 2 2 0.999843 0.999821
propublica 0 0.003287 12325 8952 87 33 0.992991 0.996327
POTUS 0 0.003139 8389 7316 0 0 1.000000 1.000000
MikeHudema 1 0.002022 12798 10474 13 8 0.998985 0.999237
AssaadRazzouk 1 0.002807 10615 7388 44 31 0.995872 0.995822
PaulEDawson 1 0.000761 10559 5675 15 8 0.998581 0.998592
ClimateBen 1 0.003454 10327 9121 135 62 0.987096 0.993248
zoenone0none 2 0.011598 24876 24855 3 2 0.999879 0.999920
studentactivism 2 0.011936 19994 19861 2 2 0.999900 0.999899
ajplus 2 0.009282 6795 5790 18 6 0.997358 0.998965
femalekissinger 2 0.001781 5197 5195 1 1 0.999808 0.999808
Refugees 3 0.009205 29470 15697 197 86 0.993360 0.994551
UN 3 0.006119 16457 12513 91 33 0.994501 0.997370
UNmigration 3 0.001623 8475 4151 278 113 0.968239 0.973499
antonioguterres 3 0.002124 6676 5626 0 0 1.000000 1.000000
nytimes 4 0.009325 21229 18971 51 16 0.997603 0.999157
guardian 4 0.002638 7924 6603 9 4 0.998865 0.999395
LeoDiCaprio 4 0.001235 5447 4897 9 9 0.998350 0.998166
NatGeo 4 0.001027 5283 4718 0 0 1.000000 1.000000
MrAhmednurAli 5 0.002175 11799 11703 1 1 0.999915 0.999915
jeremycorbyn 5 0.003088 6753 5402 3 2 0.999556 0.999630
LaylaMoran 5 0.000665 3725 3711 0 0 1.000000 1.000000
bmay 5 0.000566 3383 3029 15 13 0.995586 0.995726
BreitbartNews 6 0.000559 3076 2481 0 0 1.000000 1.000000
MaximeBernier 6 0.000966 2799 2022 17 11 0.993963 0.994589
PrisonPlanet 6 0.001274 1912 1711 1 1 0.999477 0.999416
Liz_Wheeler 6 0.000363 1810 1787 0 0 1.000000 1.000000
Kon__K 7 0.001581 9741 6362 46 24 0.995300 0.996242
bruce_haigh 7 0.000417 4512 2721 115 86 0.975146 0.969362
simonahac 7 0.000289 2775 2060 58 47 0.979527 0.977693
abcnews 7 0.000239 2051 1671 2 2 0.999026 0.998805
JustinTrudeau 8 0.003543 15733 15066 5 5 0.999682 0.999668
ianbremmer 8 0.000955 4529 4469 9 4 0.998017 0.999106
cjwerleman 8 0.000434 1052 1017 3 3 0.997156 0.997059
MusaNV18 8 0.000294 931 877 4 3 0.995722 0.996591
ndcodeine 9 0.006104 24228 24205 0 0 1.000000 1.000000
RoadtoMUT2022 9 0.000038 157 157 0 0 1.000000 1.000000
chloeprw 9 0.000016 71 71 0 0 1.000000 1.000000
themomentumco 9 0.000009 40 40 0 0 1.000000 1.000000
Table S3: Top 5 most relevant users by total inflow in the ten largest communities of the Climate Refugees dataset
User Community Pagerank Inflow In-degree Outflow Out-degree Flow ratio Degree ratio
JuanMa_Moreno 0 0.014113 37135 14986 25 8 0.999327 0.999466
alfonso_ussia 0 0.003464 17027 9717 41 30 0.997598 0.996922
okdiario 0 0.003296 13310 5792 44 11 0.996705 0.998104
AndaluciaJunta 0 0.011000 11794 6983 386 66 0.968309 0.990637
WWFespana 1 0.006797 20837 8588 1292 201 0.941615 0.977131
juralde 1 0.005168 19984 9507 363 94 0.982160 0.990209
Santi_MBarajas 1 0.003885 19234 7772 270 65 0.986157 0.991706
SEO_BirdLife 1 0.004740 10059 4557 499 138 0.952737 0.970607
FonsiLoaiza 2 0.009273 13695 9416 5 1 0.999635 0.999894
JA_DelgadoRamos 2 0.002128 10446 5413 23 12 0.997803 0.997788
Toni_Valero 2 0.002191 9947 4700 64 27 0.993607 0.994288
JavierArocaA 2 0.001303 8632 4994 84 30 0.990363 0.994029
TeresaRodr_ 3 0.002844 10018 5616 209 73 0.979564 0.987168
Ainhoasauria 3 0.003481 5677 5676 0 0 1.000000 1.000000
MiguelMorenatti 3 0.001572 4024 3916 5 4 0.998759 0.998980
lavozdelsures 3 0.000936 3343 2414 475 191 0.875589 0.926679
carlosromeroco 4 0.016908 51995 16142 4210 302 0.925096 0.981635
PartidoPACMA 4 0.004616 10101 8238 32 15 0.996842 0.998182
fromerofoto 4 0.001280 4799 1132 887 90 0.844003 0.926350
guardiacivil 4 0.004166 3612 2798 7 2 0.998066 0.999286
sanchezcastejon 5 0.032978 67372 26401 2 2 0.999970 0.999924
eldiarioes 5 0.008312 24556 11979 205 32 0.991721 0.997336
PSOE 5 0.003014 22664 7498 47 15 0.997931 0.998003
iescolar 5 0.006451 21280 10952 6 3 0.999718 0.999726
jcanadellb 6 0.000574 1478 1084 15 8 0.989953 0.992674
sninobecerra 6 0.000243 789 739 0 0 1.000000 1.000000
puntocriticoDH 6 0.000112 675 201 358 16 0.653437 0.926267
jm_clavero 6 0.000113 528 439 1 1 0.998110 0.997727
ClimateBen 7 0.002295 798 740 2 1 0.997500 0.998650
WWF_Deutschland 7 0.000356 246 199 14 5 0.946154 0.975490
iacbe 7 0.000101 173 108 13 7 0.930108 0.939130
AlanDaviesbirds 7 0.000195 158 107 22 10 0.877778 0.914530
ParlamentoAnd 8 0.002065 2855 2335 24 10 0.991664 0.995736
CiudadanosCs 8 0.000200 724 501 24 14 0.967914 0.972816
Cs_Andalucia 8 0.000254 557 440 67 32 0.892628 0.932203
AsocParqueDunar 8 0.000136 430 114 101 26 0.809793 0.814286
GotTalentES 9 0.001144 1618 1617 0 0 1.000000 1.000000
proalmerienses 9 0.000008 22 21 0 0 1.000000 1.000000
RaholaOficial 9 0.000007 22 13 16 16 0.578947 0.448276
chiquisanz 9 0.000001 10 5 308 60 0.031447 0.076923
Table S4: Top 5 most relevant users by total inflow in the ten largest communities of the Doñana dataset
S2 Language analysis
We report in Tables S5, S6, S7 and S8 the most used languages by community for the COP27, IPCC, Climate Refugees, and Doñana datasets, respectively.
Community Top 1 language Top 1 language ratio Top 2 language Top 2 language ratio
0 en 0.911 de 0.051
1 en 0.860 de 0.060
2 pt 0.980 es 0.011
3 pt 0.927 en 0.047
4 en 0.994 fr 0.001
5 fr 0.871 en 0.113
6 en 0.917 pl 0.021
7 en 0.911 fr 0.026
8 es 0.945 en 0.042
9 es 0.926 en 0.061
Table S5: Most used languages in the ten largest communities of the COP27 dataset
Community Top 1 language Top 1 language ratio Top 2 language Top 2 language ratio
0 en 0.985 de 0.003
1 en 0.773 nl 0.180
2 es 0.699 en 0.261
3 en 0.993 de 0.005
4 en 0.985 de 0.005
5 de 0.930 en 0.064
6 en 0.986 es 0.004
7 en 0.647 nl 0.323
8 en 0.791 sv 0.192
9 de 0.955 en 0.043
Table S6: Most used languages in the ten largest communities of the IPCC dataset
Community Top 1 language Top 1 language ratio Top 2 language Top 2 language ratio
0 en 0.998 ro 0.001
1 en 0.987 fi 0.003
2 en 0.990 id 0.009
3 en 0.990 es 0.003
4 en 0.991 it 0.004
5 en 0.998 it 0.001
6 en 0.990 nl 0.004
7 en 0.997 fr 0.002
8 en 0.996 it 0.001
9 th 0.991 en 0.009
Table S7: Most used languages in the ten largest communities of the Climate Refugees dataset
Community Top 1 language Top 1 language ratio Top 2 language Top 2 language ratio
0 es 0.996 pt 0.001
1 es 0.907 en 0.077
2 es 0.993 ca 0.003
3 es 0.969 pt 0.015
4 es 0.952 pt 0.024
5 es 0.991 en 0.005
6 es 0.635 ca 0.295
7 en 0.609 de 0.283
8 es 0.978 en 0.011
9 es 0.998 pt 0.001
Table S8: Most used languages in the ten largest communities of the Doñana dataset
Appendix C Additional results polarization and entropy in Twitter
In Fig. S3 we report the standarized polarization in the COP27, IPCC, Climate Refugees and Doñana networks calculated following [22] as
$\hat{\Phi}_{z}(G)=\frac{\Phi(G)-\Phi(G_{CM})}{\sqrt{<\Phi(G_{CM})^{2}>-<\Phi(G_{CM})>^{2}}}.$ (S1)
Figure S3: Standarized polarization in the Twitter dataset. Standardized measures of polarization in the retweet networks after the denoising and the calculation of z-scores.
In Fig. S4 we show the outflow entropy as a function of the outgoing isolation for the mention network between communities. We also show in Fig. S5 we show the inflow entropy as a function of the incoming isolation for the retweet network between communities.
Figure S4: Community connectivity and entropy in mention networks. Outflow entropy as a function of the normalized outgoing connectivity in the mention networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of attainable values.
Figure S5: Community connectivity and entropy in retweet networks. Inflow entropy as a function of the normalized incoming connectivity in the retweet networks of COP27, IPCC, Climate Refugees, and Doñana. The grey area corresponds to the range of attainable values.
We have computed in Fig. S6 the polarization scores in the subgraph between the two largest communities with opposing bias that speak the same language.
Figure S6: Standarized polarization in Twitter by interaction type. Standardized modularity and the E-I indices in (a) the retweet and (b) mention networks. We computed the polarization over the subnetwork of the two largest communities of opposed political bias. We have calculated the indices in the real network, in 100 realizations of the configuration model with preserved degree sequences, and standardized them.
In Fig. S7 we report the bias echo chambers in the mention network and in Fig. S8 the reliability echo chambers in the retweet network.
Figure S7: Detection of echo chambers in the mention network for political bias. Political bias of the neighborhood as a function of the bias of a user in the mention network for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
Figure S8: Detection of echo chambers in the retweet network for reliability. Media reliability of the neighborhood as a function of the bias of a user in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. studied by considering only retweets.
Figure S9: Chamber overlap matrix between users clustered using a hierarchical agglomerative algorithm. Chamber overlap between the $20$ users with more retweets in the two largest communities with opposed bias in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
The chamber overlap can distinguish between users sharing similar echo chambers (Fig.S9). We constructed the user-to-user matrices where the values correspond to their chamber overlap. Using a hierarchical agglomerative algorithm, we clustered the distance matrix and obtained an almost perfect separation of users by communities. The separation between the users in the communities is more neat in the COP27 and Doñana datasets compared to IPCC and Climate Refugees. Nevertheless, we observe a clear trend of similar echo chambers among users within the same community. In the COP27 network, the pattern is more evident for the users in the mainstream community (in blue). The separation between communities and the matching of echo chambers suggest a clearer separation in the Doñana dataset.
Figure S10: Ranking of hashtags by number of interactions in the top two opposing communities. Ranking of hashtags in the two largest communities with opposed political bias in the (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks.
In Fig. S10, we show the $20$ most tweeted and retweeted hashtags in the two communities analyzed per network. We observe a large overlap in the most tweeted hashtags between communities with a dominance of the generalistic ones.
Figure S11: Hashtag similarity in the COP27, IPCC, Climate Refugees and Doñana datasets. Similarity matrix between the set of top 20 hashtags in each of the two opposed communities. Results for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
Finally, we employ the same approach used to study the chamber overlap to characterize and compare the use of hashtags (Fig. S11). We calculated the weighted Jaccard similarity between hashtags on the fraction of tweets by community. We observe a clear grouping of the hashtags, which is much stronger in the Doñana network, indicating a stronger polarization in the use of hashtags.
Appendix D Additional results Youtube
We show in Fig. S12 the network between YouTube posts for each of the datasets analyzed.
Figure S12: Network of Youtube posts extracted from the Twitter datasets. Each node corresponds to a post and links to the number of co-commenting users. Each node is colored according to the community calculated by optimizing the modularity [75].
We have analyzed the subnetwork focusing on those communities spreading skeptical content on climate change and the larger community with mainstream content, and we report them in Fig. S13.
Figure S13: Network of YouTube posts after selecting the two main communities of mainstream and skeptics. Interaction between the posts after selecting the larger communities with mainstream and skeptical ideas. We calculated the communities by optimizing the modularity for two partitions. We identified the positioning of the communities by inspecting the videos.
Figure S14: Network of YouTube users. Networks of interactions between YouTube users. The weight of links is the number of common posts, and the color corresponds to the community calculated by the optimization of modularity [75].
Figure S15: Polarization in the network of Youtube users. Quantification of the modularity and EI indices in the user network. We have calculated the indices in the real network and in 100 randomizations where we preserve the degree sequences.