### v3

# Multifaceted polarisation and information reliability in climate change discussions on social media platforms

Abstract
Social media platforms like YouTube and Twitter play a key role in disseminating both reliable and unreliable information about climate change. This study analyses the topology of interactions in Twitter and their relation to cross-platform sharing, content discussions and emotional responses. We examined climate change discussions across four topics: the 27th United Nations Climate Change Conference, the Sixth Assessment Report of the United Nations Intergovernmental Panel on Climate Change, climate refugees, and Doñana Natural Park. While retweets reinforce in-group cohesion in the form of echo chambers, inter-group exposure is significant through mentions, suggesting that exposure to opposing views intensifies polarisation, rather than mitigates it. Ideological divides feature content differences accompanied by steeper negative sentiments, especially from right-leaning communities prone to share low-reliability information. We identified a topological alignment between platforms, indicating that ideological communities span multiple sites. Our findings show that climate change polarisation is multifaceted, involving ideological divides, structural isolation, and emotional engagement. These results suggest that effective climate policy discussions must address the emotional and identity-driven nature of public discourse and seek strategies to bridge ideological divides.
Polarisation has become a defining feature of contemporary political and social life. While a certain level of conflict and disagreement is essential to the functioning of democracies, an excessively polarised social environment can detrimentally affect their self-governing capacity [1], particularly when it pertains to issues, like climate change, that require broad consensus and coordinated action. In this paper, we examine polarisation in climate change discourse on Twitter (currently X) and YouTube, using a multidimensional approach. To capture different social and geographical contexts, we analyse four datasets related to milestone events: the 2022 UN Climate Change Conference (COP27), the release of the 6th Assessment Report by the Intergovernmental Panel on Climate Change (IPCC), the emergent issue of climate refugees, and a politically charged local debate in Spain over environmental regulation in Doñana.
The term polarisation is used across disciplines to describe the division of a population into two well-separated and opposing groups. It can be understood both as a state – the extent to which opinions or characteristics are opposed at a given time – and as a process – the movement toward greater extremes over time [2]. Furthermore, polarisation has a specific meaning related to the distribution of opinions, beliefs, attitudes, or other measurable attributes in a population (such as wealth or income) [3]. Operationalisations of this concept emphasise the formation of distinct groups in the distribution, with high homogeneity within the group and heterogeneity between groups. The core notion being that polarisation is high when a distribution is multi-modal and maximal when it is bimodal (i.e., when there are multiple distinct ’peaks’ corresponding to social groups) rather than a continuum [4].
Esteban and May [3] frames polarisation in terms of ’identification’ and ’alienation’. Polarisation is conceptualised as the result of people strongly identifying with others in their own group while feeling alienation (distance or hostility) toward those in other groups. A formal model is proposed, where society is more polarised when it contains relatively cohesive subgroups that are far apart from each other. High identification within groups means that each of the groups of individuals has internal similarity or solidarity, and high alienation means there is strong effective distance or antagonism between those groups. This idea leads to well-defined mathematical measures in which polarisation is essentially the weighted sum of pairwise ’antagonisms’ between individuals, increasing when people cluster into separated groups.
The shift of political views towards extreme positions and the formation of separate cohesive groups often is related to the presence of echo chambers, where interactions predominantly occur among individuals who share similar beliefs and ideas [5, 6, 7]. As interest in measuring [6, 8, 9, 10, 11] and modelling [12] the dynamics of polarisation and echo chambers grows, efforts to identify the factors influencing the interaction patterns face significant challenges. The recommendation algorithms employed by these platforms influence these dynamics [13, 14, 15] by limiting exposure to opposing viewpoints in users’ feeds [16]. Yet the rise of polarisation is not only an effect of homophily as it has been shown that individuals are often exposed to opposed opinions [17].
While polarisation is a natural aspect in pluralistic societies, its intensification over the last decades is a matter of debate both among academics and in public discourse [18, 2, 19, 20]. Multiple studies point to the fact that the widespread adoption of social media has driven an increase in polarisation as observed across experimental settings [21] and various platforms [8, 16, 22]. However, its presence and magnitude vary across subjects, can be triggered by external events [23, 24, 25, 26] and is steeper around political discussions than in more general subjects [27]. Platforms also play a role, while polarisation has been observed on Twitter, YouTube, or Instagram [28, 29, 30, 31], in others, such as WhatsApp, there is a depolarisation effect [32]. Even though the interest in measuring social and political polarisation dates from before the digital era [33], the widespread use of online social media and data accessibility sparked the development of frameworks to quantify it. The techniques developed vary depending on the perspective, while the topology of interactions provides insights into the preferences to form social bonds based on political views [34], textual analysis can capture the divergence in emotions and policies [35]
The combination of a polarised environment [36] and behavioural factors, such as confirmation bias [37, 38], has been closely associated with the dissemination of low-reliability information [39], influencing the outcome of presidential elections [40] and increasing vaccine hesitancy [41]. Despite these challenges, there have been concerted efforts to combat the spread of fake news [42]. Nonetheless, debates persist regarding the distinction between reliable and unreliable news cascades. Although initial studies highlighted structural differences [43], subsequent analyses have suggested that disparities may be more attributable to differences in cascade size [44].
Polarisation manifests across political issues, but also other kinds of events and subjects as we have witnessed an increase in climate change discussions and related events [45, 46, 20, 26]. Previous research has focused on the origins of this polarisation [47], its relation to political affiliations [48], and the influential roles played by media outlets [49] and corporations [50]. Recently, an increasing polarisation has been reported in locations affected by extreme weather events or environmental disputes [46]. The sentiment towards climate change mitigation measures varies significantly across regions, shaped by local political and cultural contexts [51]. While the spread of low-reliability information in these contexts has been noted, its connection with polarised environments remains unclear [52, 53, 54]. Our study addresses this gap by studying a highly localised and politically charged debate over environmental protection in Spain, Doñana, where the interplay between misinformation, partisanity, and institutional trust is especially salient.
The broad concept of polarisation can then be further specified, with particular relevance to our work, by focusing on the notions of political and affective polarisation. Political polarisation usually refers to the distance between political preferences, identities, or parties within a society [2]. In relation to this, the analysis of a retweet network relies on the empirical and theoretically supported assumption that the way users share information is related to and a source of their collective political identity [55, 56]. In the same line, information sharing is related to the practice of networked gatekeeping that put in relation elite and mass polarisation [57] and lead to the formation of affective publics [58]. Affective polarisation is also emphasised in political science and is distinct from policy disagreement. It describes the phenomenon of partisans feeling not just different, but actively antagonistic toward each other. [19, 59]. In other words, it’s the polarisation of feelings or attitudes toward the social groups representing the other, rather than toward policy issues themselves. This is well captured through expressions of antagonism between groups, such as emotional tone in communication. In relation with the structural polarisation metrics we apply in this work, it is central to the notion of how users use retweet and mention as different ways of signalling in- and out-group relations and identities [60].
The initial metrics to quantify the macroscopic level of structural polarisation primarily concentrate on modularity and the analysis of community boundaries [61, 62, 34]. However, those only provide a partial perspective on the organisation of social connections, as they largely disregard the multi-scalar organisation of networks and the role of individual nodes. Metrics designed to capture long-range structural properties include label propagation [63], centrality measures, and random walk diffusion [64]. Even though most assume a dipole organisation, recent developments can capture multipole polarisation [65]. Structural measures in dealing with the networked organisation of social relations [66] provide a measure of polarisation being content-agnostic. Natural language processing techniques are essential for assessing the content of the divergence on policy-related issues [32, 45] and evaluating the emotions of individuals and their interactions [60]. While our approach focuses on structural and affective dynamics at a group level, recent studies have aimed to integrate these with individual-level opinions [31] providing further context into the structural organisation of different ideological perspectives.
Building on previous literature that tends to isolate structural, ideological, or affective polarisation, this study offers an integrated approach to capture their intersection in climate change discourse. Drawing on Twitter and YouTube data, we examine how community structure, emotional tone, and media reliability interact across platforms and contexts. To do so, we combine network analysis (to assess structural patterns), ideological measures (using political bias and source reliability), and sentiment metrics (focused on toxicity in quote tweets and cross-group mentions). This multidimensional approach allows us to observe how different dimensions of polarisation reinforce one another. Users with similar ideological leanings tend to cluster in retweet networks and express aligned affective reactions, particularly in politically salient debates. Interactions across communities, while rare in the form of retweets, are more frequent through mentions and often carry antagonistic sentiment—suggesting emotional confrontation despite low structural connectivity.
Our main contribution is the development of an integrative framework that captures structural, ideological, and affective polarisation across digital platforms. Applying this framework to four climate-related events, we examine how polarisation manifests differently depending on the topic and geographical context, ranging from global summits to locally contentious environmental debates. This multidimensional analysis reveals consistent links between the structure of online communities, the emotional tone of user interactions, and the reliability of the content being shared. Notably, we observe that polarisation is not confined to a single platform: ideologically aligned groups on Twitter also engage with similar content on YouTube, indicating a broader cross-platform coherence. These findings underscore the importance of moving beyond single-platform or single-dimension studies in order to understand the full complexity of polarisation in digital climate discourse.
I Data Description
We extracted four datasets related to climate change topics from Twitter through the Twitter Academic API. These datasets correspond to the 2022 United Nations Climate Change Conference (COP27), the release of the 6th Assessment Report by the Intergovernmental Panel on Climate Change (IPCC), the climate refugee crisis resulting from extreme weather conditions (Climate Refugees), and to legislation affecting the natural reserve of Doñana in the south of Spain (Doñana). Tweets were collected using Twarc [67] by searching for keywords related to each discussion. The keywords selected were closely related to each event. For COP27 the keywords included COP27 and related concepts such as LossAndDamage. For IPCC the keywords were IPCC and its official account (@IPCC_CH). For climate refugees the keywords were related combinations included climate refugee(s) and climate migration. For Doñana only the keyword Doñana was used. The datasets for specific events (COP27 and IPCC) were filtered within the span of the pre and post-event dates. The Climate Refugees and Doñana datasets span a longer period compared to the IPCC and COP27 datasets. The IPCC dataset, in particular, covers only nine days, coinciding with the peak discussion surrounding the release of the 6th Assessment Report (See Table 1). We extracted the corresponding YouTube dataset by analysing the URLs referenced in each Twitter dataset and downloading the posts and comments using the YouTube Data API. To ensure relevance, we filtered the posts by searching for occurrences of specific keywords in each video title or description (Section S1).
Min. date Max. date Users Tweets Retweets Replies Quotes
COP27 2022-09-01 2022-11-27 1,351,903 866,753 4,977,874 205,973 175,078
IPCC 2023-03-18 2023-03-26 157,056 31,138 267,971 45,863 7,751
C. Refugees 2008-03-10 2022-12-31 841,454 384,267 1,376,057 139,699 38,909
Doñana 2019-01-01 2023-04-30 290,782 139,478 1,187,646 135,245 25,056
Table 1: Summary table of the Twitter datasets analysed. Minimum and maximum dates, observed number of users, and number of tweets by typology.
Videos Min. Date Max. Date Avg. views Avg. likes Avg. comments
COP27 624 2008-02-11 2023-05-12 571,964.35 5,860.32 306.99
IPCC 145 2007-11-03 2023-06-09 350,178.66 8,991.67 1,455.16
C. Refugees 215 2009-11-22 2023-11-06 243,720.27 3,604.81 527.40
Doñana 191 2009-01-11 2023-06-01 101,098.32 3,622.68 246.25
Table 2: Summary table of the YouTube datasets analysed. Number of videos in each community, minimum and maximum dates observed, and average number of views, likes, and comments.
Unlike the Twitter dataset, the period of the comments is longer due to users frequently referencing older videos. Each dataset exhibits peaks at different times based on its unique characteristics. In particular, we observe a strong alignment in the IPCC and COP27 datasets with the event dates (Table 2). Additional insights, including the data collection procedure can be found in Section S1 of the Supplementary Material.
II Results
II.1 Network analysis and structural polarisation
We evaluate the structural polarisation on Twitter by considering the unweighted network of retweets with undirected links indicating a retweet interaction between two users. Previous studies have used retweets as a proxy for the influence between individuals allowing for the computation of their latent ideology [63, 45]. Similarly, we aim to measure how the influence between individuals shapes network organisation. We applied the Metis partition algorithm [68] to split the graph into two clusters, and we calculated the modularity [69], the E-I index [70] and the adaptive E-I index [71] that adjusts for the cluster sizes. These metrics evaluate the topological interactions between users, the mesoscale structure and the extent to which users in different groups interact with one another without considering the content. The results for the four datasets analysed are presented in Fig. 1. Following [34], we computed polarisation metrics in the observed network $\Phi(G)$, as well as in the ensemble of configuration models with equivalent degree distribution $\Phi(G_{CM})$ [72], and calculated the denoised value $\hat{\Phi}(G)=\Phi(G)-\Phi(G_{CM})$. For completeness, we report the standardised values [34] in Fig. S3.
Figure 1: Polarization in the Twitter dataset. Quantification of the polarization in Twitter: (a) modularity, (b) E-I index, and (c) adaptive E-I index. We provide the measures for the observed networks in red, across 100 realisations of the configuration model with preserved degree sequences in blue, and the difference between them (denoised value) in yellow.
All networks exhibit structural polarisation according to the denoised value, suggesting the presence of two groups of strongly connected individuals. In particular, the COP27 and Doñana datasets displayed larger denoised values, suggesting a stronger polarisation around the topics. The higher structural polarisation observed in the COP27 dataset may be attributed to transnational spheres using different languages that do not interact with each other [73]. The first and second largest communities, in terms of the number of messages, are dominated by English. The third and fourth communities primarily consist of Portuguese messages, followed by others where French and Spanish are more prevalent. Further details on this distribution per dataset and community can be found in Section S2.2. The Doñana dataset captures the discussion around a new law related to natural park preservation, which involves political parties and users with strong political affiliations like politicians, journalists, or activists. The lower polarisation observed in the IPCC dataset may be due to its nature as a scientific report release rather than a mainstream international conference like COP, potentially attracting a more scientifically educated audience and inducing a lower level of polarisation. Although the largest community (labelled as 0) is composed mainly of scientists, primarily from the UK, other communities consist of activists (community 3) or politicians (community 6), such as @GretaThunberg and @BarackObama, respectively.
Figure 2: Community interaction network. Interaction between the ten largest communities in (a, b) COP27, (c, d) IPCC, (e, f) Climate Refugees, (g, h) Doñana. We report the interactions via (a, c, e, g) retweets and (b, d, f, h) mentions. Arrows indicate the direction of the interaction, the colour intensity and width, the volume. The size of the dots represents community size. In blue, we show the interaction through retweets and, in orange, through mentions.
While the retweet network has been previously interpreted as the influence between social media users, other types of interactions, such as mentions, often feature different structural properties [61]. We examine the network of retweets and mentions among the $10$ largest communities of the networks in Fig. 2. We identified these communities using modularity optimisation on the directed weighted graph of retweets only [69], and then we incorporate mentions. This approach allows us to identify clusters based on influence and later analyse how they interact through mentions and retweets. The interaction patterns among the largest community (labelled as 0) and the remaining communities vary across datasets and types of interactions. In the COP27 and IPCC networks, the smaller communities mostly retweet content from the largest community, which is more isolated and engages in fewer interactions with the rest of the network. The largest communities include official accounts and major figures, such as the Secretary-General of the United Nations (@antonioguterres) or the Intergovernmental Panel on Climate Change (@IPCC_CH), representing COP27 and the IPCC, respectively. Conversely, the Doñana network exhibits higher flow heterogeneity between communities, with community 5 playing a central role in attracting most of the flow. The detailed overview of the users in the community reveals the presence of the president of Spain (@sanchezcastejon), the ruling political party (@PSOE), and sympathiser journalists (@eldiarioes, @iescolar). There are significant structural differences depending on the type of interaction. For example, in the IPCC network, there are relatively few retweets from community 1, with accounts concentrated flow such as @uksciencechiefUK (Government Office for Science) and @BernieSpofforth (the anti-climate movement figure) to community 0, formed by well-know climate science accounts such as @ed_hawkins or @MrMatthewTodd. However, there is a high volume of mentions between these communities (see details in Section S2.1).
Figure 3: Reliability as a function of political bias in the studied networks. Density plot showing the distribution of reliability as a function of bias for the communities detected in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. For each dataset, the two communities with the largest and most divergent political views are identified. In dark brown the left-biased community and in light brown the right-biased community
II.1.1 Isolation and entropy
We further evaluate the connectivity patterns between communities by quantifying the structural properties of their inter- and intra-group interactions. To quantify the isolation of communities, we calculated the directed interactions between them, and we computed the normalised outflow difference given by
$\Delta F^{out}_{i}=\frac{\sum_{j,j\neq i}T_{ij}-T_{ii}}{\sum_{\forall j}T_{ij}},$ (1)
where $T_{ij}$ is the number of interactions from community $i$ to community $j$. $\Delta F^{out}_{i}$ provides a measure of a community’s isolation. A value of 1-1 indicates that a community has only internal flow, whereas a value close to $1$ suggests that most of the flow is directed toward other communities. A value of 0 signifies that half of the flow is external to the community, while the other half remains internal.
To understand the diversity of interactions among communities, we compute the entropy of flows as
$H^{out}_{i}=-\sum_{\forall j}p_{ij}logp_{ij},$ (2)
where $p_{ij}$ is the probability of having an interaction from $i$ to $j$ given by $T_{ij}/\sum_{\forall j}T_{ij}$. While $\Delta F^{out}_{i}$ quantifies the ratio between external and internal links and goes from 1-1 (when all flows are internal) to $1$ (when all flows are external), the entropy ($H^{out}_{i}$) provides information on the variety of those flows. The entropy is 0 when there is no variety of flows and $1$ when it is maximum. Most of the largest communities exhibit a certain degree of isolation, indicated by negative values. In the COP27 network, the most isolated communities are 2 and 3, both Brazilian ($\Delta F^{out}_{2}=-0.97$ and $\Delta F^{out}_{3}=-0.94$), followed by community 1 ($\Delta F^{out}_{1}=-0.88$), which is the most isolated and has the lowest entropy ($H^{out}_{1}=0.05$) among English-speaking groups. This community consists of sceptics with influential accounts such as @JamesMelville and @BernieSpofforth. Following the same pattern, in the IPCC network, community 1 has the highest isolation value ($\Delta F^{out}_{1}=-0.98$) and the lowest entropy ($H^{out}_{1}=0.014$). This community also includes, among other accounts, @BernieSpofforth. Similarly, in the Climate Refugees network, community 6 is the most isolated ($\Delta F^{out}_{6}=-0.93$) and has the lowest entropy ($H^{out}_{6}=0.03$), featuring accounts such as @MaximeBernier and @PrisonPlanet. In the Spanish Doñana network, community 0 is the most isolated ($\Delta F^{out}_{0}=-0.95$) and has the lowest entropy ($H^{out}_{0}=0.02$). Across all datasets, communities with right-wing alignment demonstrate strong isolation and low entropy, implying that their messages have limited spread across the network.
However, the pattern changes when considering the mention graph. The largest mainstream communities show far less mentions of other communities, for instance, community 0 on IPCC and COP27 networks has high isolation ($\Delta F^{out}_{0}=-0.5$ and $\Delta F^{out}_{0}=-0.57$, respectively). Whereas the low-reliability communities display lower isolation and more external mentions, its the case of community 1 in COP27 ($\Delta F^{out}_{1}=0.37$ and $H^{out}_{1}=0.25$). These results reinforce, first, the heterogeneity in group connectivity depending on the interaction types and, second, different levels of isolation by communities.
II.2 Political polarisation
II.2.1 Bias and reliability
We assess if the structural polarisation is related to the divergence in political views and information diet by calculating the bias and reliability of the URLs shared by the users. Denser connections between users with similar ideological biases could indicate a closer influence between them and a greater disagreement on climate change discussions. Information on the political bias and reliability of news outlets was obtained from MediaBias fact-check [74], where positive and negative political bias correspond to right-wing and left-wing media, respectively. Outlets with higher reliability are considered more trustworthy and tend to share more verified news. For the Doñana dataset, the reliability and bias were based on a separate dataset [75], given that most media sources are Spanish. To facilitate the analysis, we have normalised the bias within the range [-1, 1] and the reliability within the range [0, 1]. The minimum and maximum biases are set according to the higher absolute value observed. In Fig. 3(a)-(d), we show the density plot of community reliability as a function of their political bias according to the shared URLs in the IPCC, COP27, Climate Refugees, and Doñana networks. We computed user-level values by averaging over the sources they shared and then calculated community-level values by averaging over the users. All datasets exhibit a left-wing bias, as most communities show negative values, especially in the IPCC and Climate Refugee networks. Possible explanations include a systemic bias in media coverage of these topics, the tendency of climate change sceptics to share information from unconventional sources that are not explicitly flagged by media bias indicators, or an over-representation of left-wing users in climate change discussions, as they may be more likely to engage due to greater awareness of the issue. The left-wing bias is particularly pronounced in the IPCC and Climate Refugee networks, where the cluster of large communities is centred around negative values. However, the bias is more dispersed in the Doñana dataset, possibly due to differences in the shared media sources or a wider ideological variety.
The communities with right-wing bias we have identified are: community 1 in COP27 (with key actors as @JamesMelville) and IPCC (with the presence of @BernieSpofforth), community 6 in Climate Refugees (with @MaximeBernier or @PrisonPlanet), and community 0 in Doñana (with @JuanMa_Moreno or @alfonso_ussia). We performed significance tests to compare the bias and reliability of communities, revealing several significant differences. In terms of bias, the most notable differences (p-value ¡ 0.001) are observed in the previously mentioned right-wing biased communities compared to the other communities. Regarding reliability, community 4 in COP27 (with main actors such as @CarolineLucas), community 3 in IPCC (with @GretaThunberg or @CarolineLucas), community 5 in Climate Refugees (@jeremycorbyn), and community 1 in Doñana (@WWFespana and other conservation organisations) show the most significant higher reliability relative to the other communities (see Figs. S6-S7).
Based on the previous findings and manual inspection of the communities, we identified the two largest communities with opposing political leanings. In the IPCC and COP27 datasets, community 0 represents the largest community aligned with the mainstream climate change narrative, while community 1 is the primary advocate of a counter-narrative message. However, in the Doñana network, the labels are swapped, as the right-wing biased community has a larger user base. For the Climate Refugees network, most of the top 10 communities exhibit similar bias-reliability values, except for community 6, which stands out with low reliability and a right-wing bias. In this dataset, we designate community 1 as the primary left-wing biased community, as it includes well-known climate change activists such as @MikeHudema or @GretaThunberg and engages in more interactions with the low-reliability community. Our results point out that most communities in climate change discussions share reliable information and have a left-biased alignment. However, there is at least one large community in each dataset on the right-side alignment that shares less reliable information.
These results indicate differences in the network topology depending on the type of interactions. Retweets, which are linked to influence, connect individuals in communities with similar biases. Instead, a higher number of mentions across communities suggests interactions between individuals from opposing sides of the ideological spectrum.
Given that many networks include users from different countries speaking various languages, we also measured polarisation by focusing on the two largest communities that speak the same language but have opposing biases, comparing the retweet and mention networks. The results still indicate polarisation in the retweet network, with modularity values ranging from $0.07$ in the IPCC network to $0.21$ in the Doñana network. Even in the mention networks—where structural scores are lower (modularity is $0.05$ in COP27, $0.16$ in Climate Refugees, and $0.15$ in Doñana)—only the IPCC network shows no polarisation (modularity = 0.23-0.23). These findings further support the observed disparity in network organisation based on interaction typology (see Section S2.3 for more details).
II.2.2 Echo chambers
Throughout this section, we assess the presence of echo chambers with a focus on the two largest communities exhibiting opposing political biases. Our objective is to assess the interaction patterns between users as a function of their bias. We employed an approach developed in [9] to evaluate the existence of echo chambers. For a given user $i$ with degree $k_{i}$, we computed its average bias $x_{i}$ and reliability $y_{i}$, along with the average bias and reliability of its neighbourhood, denoted as $\frac{1}{k_{i}}\sum_{\forall j\neq i}A_{ij}x_{j}$ and $\frac{1}{k_{i}}\sum_{\forall j\neq i}A_{ij}y_{j}$, respectively.
Figure 4: Detection of echo chambers in ideological biases for the retweet networks. Neighbourhood political bias as a function of the user bias in the retweet network for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
We show the density plot for the bias of the neighbourhood of users as a function of their values when considering the interactions through retweets in Fig.4. As expected from the overall ideological positioning of the users, there is a hotspot on the left of the ideological spectrum and a weaker one on the right side. We have identified a diagonal trend that could be related to the presence of echo chambers. Right-wing biased users interact more with left-wing biased users, as there are two hotspots for positive values in some cases. The presence of echo chambers is more evident in the Doñana dataset than in the rest of the networks, mainly due to the presence of more users with right-leaning discourse. There are three hotspots of comparable magnitude across the diagonal with right-, centre-, and left-leaning discourses. Still, the hotspot on the left side of the spectrum displays a lower ideological isolation, since it interacts with users featuring a bias close to 0. The right-wing biased users interact mainly with each other and not with users in the centre. When analysing other types of interactions, such as the mentions, the presence of echo chambers is weaker with nearly horizontal hotspots in the COP27 and IPCC datasets (see details in Figs. S11). This observation is consistent with the results on polarisation from the previous section. The existence of echo chambers is less pronounced when considering different types of interactions, and mentions could serve as a bridge between users with different political biases.
We also examine the echo chambers based on the reliability of the content shared by users and their neighbourhoods, where we find three hotspots for high, medium, and small values. The hotspots in high reliability suggest that users who share highly reliable sources interact with each other, while low-reliability users interact with users who spread highly reliable sources (see Fig. S12).
We inspect in detail the user interaction similarity by quantifying their echo chamber [76], without considering the ideology. Given a set of leading users $i$, which have an audience $A_{i}$ composed of the set of users that retweeted it, their chamber $C_{i}$ is the set of users retweeted by audience $A_{i}$. We focus on the two communities with opposing political biases, specifically the top twenty users with the most retweets in each of those communities. The first quantity we will analyse is the chamber overlap between a user $i$ and a user $j$ given by
$q_{ij}=\frac{C_{i}\cap C_{j}}{C_{i}\cup C_{j}},$ (3)
where $C_{i}$ and $C_{j}$ are the chambers of users $i$ and $j$. It is worth noting that the quantity $q_{ij}$ is symmetric by definition and is equal to $1$ when chambers are identical and 0 when they are completely different.
We compute the chamber overlap distribution $P(q_{ij})$, where a notable disparity between networks is observed (see Fig.S13). A strong bimodal distribution suggests the presence of echo chambers, as each user has low values of $q_{ij}$ with users outside its echo chamber and high values with those inside it. The results from the COP27 and Climate Refugees datasets suggest that leaders have more heterogeneous chambers compared to the IPCC and Doñana datasets, where we observe a peak for high chamber overlap. The Doñana dataset features a strong echo chamber effect with two separated peaks at low and high values of $q_{ij}$. We have performed hierarchical clustering of the top 20 users per community based on the similarity of their chambers, showing a strong alignment between the communities detected (for more details see Section S2.6).
Figure 5: Hashtag polarisation. Polarisation around hashtags for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. Negative and positive values correspond to hashtags more abundant in the left-biased and right-biased communities, respectively.
II.3 Policy and affective polarisation
On account of the structural polarisation observed, with groups of users sharing political bias, we further investigate if there are perceptible differences in their climate change views. Structural metrics are purely based on user interactions and, thus, provide limited information on the reasons and motivations of the groups. To get a first assessment of the content discussed in each of the two opposed communities and evaluate if they have confronted perspectives on climate change, we inspect the patterns of hashtags shared. Despite their ideological confrontation, they share most of them, likely due to the common subject of discussion (Fig. S15). The mainstream communities use rather generalist hashtags such as #COP27\#COP27, #IPCC\#IPCC, #climatechange\#climatechange, #ar6\#ar6, or #climatereport\#climatereport. In Doñana, common hashtags between communities are less frequent, which could indicate a stronger political polarisation. We observe hashtags related to the environment as #donana\#donana, #donanaseextingue\#donanaseextingue, or #salvemosdonana\#salvemosdonana for the mainstream community and hashtags related to right-wing parties as #teamvox\#teamvox, or against the left-wing government as #gobiernodimision\#gobiernodimision, in the right-wing biased community.
To compute the hashtag polarisation, we calculated the fraction of tweets with each hashtag by community, given by
$P_{i}^{\#}=\frac{S^{l}N_{i}^{r}-S^{r}N_{i}^{l}}{S^{r}N^{l}_{i}+S^{l}N_{i}^{r}},$ (4)
where $N_{i}^{l}$ and $N_{i}^{r}$ are the number of tweets with the hashtag $i$ in the main left- and right-wing communities, respectively, and $S^{l}$ and $S^{r}$ the number of users in each community. This formulation takes into account the uneven sizes of the communities. We display the polarisation of the top hashtags in Fig. 5. The number of hashtags in each network varies since we have considered the $20$ most shared in each community, and some of them can overlap. Interestingly, we observe only a few hashtags with values close to 0, suggesting a strong polarisation in the use of hashtags. In the IPCC and COP27 datasets, the mainstream community dominates most of the hashtags, while the community spreading a counter-narrative dominates only a few hashtags related to the denial of climate change that include #climatehoax\#climatehoax or #agenda2030\#agenda2030. In COP27, we also observe hashtags related to environmental hypocrisy such as #greenwash\#greenwash. In the Doñana network, we also observe only a few hashtags close to zero and hashtags almost equally divided between +1+1 and 1-1. We performed a Z-test to assess the significance of hashtag frequency differences between the most polarised communities. Among the hashtags with highest absolute Z-scores, #climatescam\#climatescam is predominantly used by right-leaning communities in both the COP27 and IPCC networks. In the Climate Refugees network, one of the hashtags with a high Z-score is #howdareyou\#howdareyou, again primarily used by the right-wing community. In the case of the Spanish discussion about Doñana, the hashtags with the highest Z-scores are those criticising the Spanish government, as previously mentioned (see Section S3.1 for details).
Whereas hashtags can provide a glimpse of the ideas discussed within each community, they offer a limited perspective to the opinion of individuals as they can be used in different contexts. To better understand the divergence in climate change views, we conducted a topic analysis on tweets from the two largest communities with opposing political perspectives. We assess topic polarisation by quantifying the number of users within each community who discuss a given topic, using Eq. 4. Most topics are either dominated by the left-biased community or the low-reliability community. In the COP27 network, the right-leaning discourse reveals corruption discussions with money-related terms, and political references to the UK, China and India. In contrast, the left-biased community focuses on essential resources such as water and food, as well as the urgency of taking action. In the IPCC network, topics within the right-biased community express scepticism about the scientific basis of climate change predictions, reference COVID-19, and include discussions that are often unrelated or misleading (for more details about the topics see Section S3.2). The results confirm that it could be composed of individuals posting low-reliability information. The left-biased community emphasises the need to reduce emissions and the urgency of taking action, using language similar to that found in COP27 discussions. In the Climate Refugees network, the right-biased community focuses on jobs and border security, whereas the left-biased community highlights the risks posed by natural disasters. The greatest divergence in discussions is observed in the Doñana network. The right-biased community focuses on the political handling of the situation, with an emphasis on the Spanish president, the ruling party (PSOE), and the European Union. In contrast, the mainstream community centres its discussion on the Doñana National Park and its conservation.
Figure 6: Average negative sentiment in messages in community interactions. Average negative sentiment of the messages as a function of the poster community (y-axis) and quoted (x-axis). The results correspond to (a) COP27, (b) IPCC,(c) Climate Refugees, (d) Doñana.
Besides the disparity in the content shared, via hashtags and topics, we want to adress whether those groups are aware of each other and if they are emotionally aligned. Therefore, We evaluate emotional repulsion in relation to political views by analysing the sentiment of quote messages. Specifically, we assess the average negative sentiment based on the community of both the sender and the receiver (Fig. 6). We have implemented a sentiment analysis on the four Twitter datasets using a RoBERTa-based model fine-tuned on around 58 million of tweets [77, 78]. Overall, messages from right-wing biased communities exhibit a higher degree of negative sentiment, particularly when directed at left-wing biased communities (0.48 for COP27, 0.54 for IPCC, and 0.55 for Climate Refugees). The difference between intra- and inter-community messages is most pronounced in the IPCC and Doñana networks. In the IPCC network, the negativity score for right-leaning to left-leaning interactions is 0.54, while for right-leaning to right-leaning interactions, it is 0.46 ($\Delta=0.08$). Similarly, in the Doñana network, right-leaning to left-leaning interactions have a negativity score of 0.38, compared to 0.31 for right-leaning to right-leaning interactions ($\Delta=0.07$). Interestingly, the most pronounced difference in sentiment between intra- and inter-community interactions occurs among users in the left-wing biased community. In the extreme case of the COP27 network, left-leaning users display a $\Delta=0.23$ difference, while also being associated with the dissemination of more reliable information. Across all networks, the lowest levels of negative sentiment are observed within the left-wing biased communities, ranging from 0.20 in COP27 to 0.39 in Climate Refugees. In contrast, the right-wing biased community exhibits consistently higher negativity, even within its own interactions, with values ranging from 0.31 in Doñana to 0.53 in Climate Refugees.
We also conducted a Mann-Whitney U test to evaluate the statistical significance of negativity differences between communities. A consistent pattern emerges in the COP27, IPCC, and Climate Refugees networks: right-leaning to left-leaning interactions display significantly greater negativity than all other combinations. For example, in the IPCC network, the negativity of right-leaning to left-leaning interactions is significantly greater than that of left-wing biased to right-wing biased interactions ($p<.001$), left-wing biased to left-wing biased interactions ($p<.001$), and right-wing biased to right-wing biased interactions ($p<.001$) (see Fig. S19 for additional comparisons). Furthermore, across all datasets, right-wing biased to right-wing biased interactions exhibit significantly higher negativity compared to left-wing biased to left-wing biased interactions (COP27: $p<.001$; IPCC: $p<.001$; Climate Refugees: $p<.001$; and Doñana: $p<.001$). A similar pattern is observed when comparing left-wing biased to right-wing biased interactions against left-wing biased to left-wing biased interactions (COP27: $p<.001$; IPCC: $p<.001$; Climate Refugees: $p<.001$; and Doñana: $p<.001$). Hence, the structural communities identified are aligned with a divergence on content discussion related to climate change and stronger emotional responses. The results suggest that polarisation on the subject takes multiple forms.
II.4 Cross-platform analysis
We have analysed the dataset created from the YouTube videos quoted and referenced on Twitter to assess if the observed polarisation facets are aligned with content consumption in video platforms as partisan media affects the dynamics of social networks [79]. The information gathered for each post includes the transcription and description of each video, the corresponding channel, and the user comments. The comment information allows us to create a bipartite network between users and posts since we know the users who have commented on each video. We can project the bipartite network into a post network or a user network. In the post network, the weight corresponds to the number of users in common and, in the user network, to the posts in common. The graphs are undirected by construction in both cases. We performed a community analysis using the greedy modularity optimisation on the weighted network [69]. The organisation of communities do not display a clear separation between communities with several interconnections except for the Doñana network, where communities zero and one have a clear separation (see Section S4).
Figure 7: Interaction between Twitter and YouTube communities. Number of references to YouTube communities by the top six Twitter communities in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. To account for the uneven size of YouTube communities, we have divided the counts by the YouTube community size.
We conducted an analysis linking Twitter and YouTube social media platforms to study the cross-platforms effects. We create the Twitter subnetwork of users who shared YouTube links related to climate change, tagged according to the YouTube community of the videos. The connection between network organisation and the shared YouTube content is clearer in the Doñana and IPCC networks. In both cases, users tweeting videos from YouTube community zero are more tightly connected than other users (see the subnetwork in Figs. S22 and S23).
Furthermore, we computed the frequency with which each Twitter community references YouTube posts and their corresponding communities. To account for the heterogeneity in community sizes observed in the YouTube dataset, we normalised the number of references to each YouTube community by its size. A strong connection exists between Twitter and YouTube communities, as most references within each Twitter community align with a specific YouTube community. Additionally, each Twitter community exhibits a unique pattern of references to YouTube.
In the COP27 network, the larger left-wing biased Twitter community 0 predominantly references YouTube community 0 ($35.8\%$), while the larger right-wing biased and low-reliability community 1 primarily references posts in YouTube community 2 ($57.3\%$). Similarly, in the Climate Refugees network, Twitter community 6, known for spreading low-reliability information, interacts significantly with YouTube community 2 ($74.6\%$). Finally, in the Doñana network, community 0 from both social networks interacts intensively ($84.3\%$). Conversely, in the IPCC network, the larger right-wing biased community 1 references multiple YouTube communities, specifically communities 1 ($20.2\%$), 2 ($35.3\%$), and 5 ($23.4\%$), suggesting a tendency to share diverse typologies of YouTube content. With the exception of the IPCC network, posts shared by the larger right-wing biased community on Twitter predominantly align with a single YouTube community (Fig. 7).
III Discussion
Our results confirm that polarisation in climate change discussions manifests in different forms through structural, ideological, and emotional components. We also observed distinct views on climate change and patterns of content consumption. The structural polarisation around the subject is aligned with distinct political biases and information reliability consumption. Not only that, the topological communities also show an alignment on perspective towards the environmental crisis and a stronger emotional response. Overall, polarisation takes multiple aligned forms. While previous work often emphasised the role of ideological segregation, our cross-platform analysis shows that inter-community exposure remains significant—especially via mentions—but is frequently accompanied by affective repulsion and asymmetric negativity, particularly from low-reliability, right-leaning communities.
This suggests that polarisation is not merely a consequence of echo chambers or algorithmic filtering, but also reflects active processes of antagonism and identity signalling, in line with theories of affective polarisation. Even in multilingual and transnational contexts, such as the COP27 dataset, we observe stable polarisation structures shaped more by political affiliations and emotional tone than by language or geography. Moreover, the structural signature of interactions varies across platforms and interaction types: while retweets act as endorsement signals and tend to reinforce in-group cohesion, mentions frequently bridge ideological divides, albeit often with antagonistic or negative sentiment.
The interplay between information reliability and political alignment also emerges as a key factor. Right-leaning communities consistently share lower-reliability content and exhibit higher isolation and negativity, yet they remain engaged with mainstream communities—albeit often in confrontational terms. This pattern challenges the notion that low-reliability information spreads in isolation and points instead to a contested media space where political identity, trust, and emotion are tightly intertwined. Our results align with other studies showing that, despite higher engagement within ideologically similar groups, users are still exposed to opposing views [17]. However, such exposure does not necessarily mitigate polarisation. On the contrary, it has been shown that confrontation with opposing views can intensify polarisation instead of reducing it [21]. This is reflected in our sentiment analysis, where messages from right-biased communities consistently show higher levels of negativity—especially when directed toward left-leaning users [60]. Overall, the sentiment in the low-reliability community was more negative. Our results are in line with recent research [26], which finds that climate change dissenters are closely related and tend to engage with a left-biased mainstream community [26, 60, 80]. Furthermore, toxicity levels are not evenly distributed across interaction types: replies and mentions are significantly more toxic than other forms of engagement [20]. This observation suggests that low-reliability individuals may reference the mainstream community to engage in discussions but get little attention in return [20], highlighting an asymmetry between left- and right-leaning communities [81].
The cross-platform alignment between Twitter and YouTube reinforces the idea that ideological communities are transversal across platforms, with users not only engaging with like-minded individuals but also selecting and amplifying content from ideologically aligned sources elsewhere. This calls for a broader, integrated approach to studying online polarisation, moving beyond single-platform analyses. Additionally, our results indicate that the degree and manifestation of polarisation can vary substantially across topics and events—suggesting that it is not static, but shaped by contextual factors such as the political salience of the issue or the type of event. This may reflect a dynamic of partisan sorting, whereby individuals increasingly align their social media behaviour with broader political identities, as already observed in the Finnish Twittersphere [71]. Prior studies, such as [54], observed that the 2019 IPCC Special Report on Climate Change and Land sparked significant public discourse on social media, with dietary recommendations becoming a highly polarised topic. Similar to our findings on localised polarisation patterns (e.g., Doñana), that report showed that climate solutions affecting lifestyle choices may trigger elevated levels of contention on platforms like Twitter. The alignment between platforms could be amplified by recommendation algorithms, which can function as filter bubbles that reinforce users’ beliefs [13, 82]. However, recent studies have downplayed the contribution of algorithms to polarisation [83]. Our results support the idea that polarisation is ubiquitous across platforms, but not driven solely by lack of exposure to opposing viewpoints.
These findings have important implications for the broader climate change debate. The persistence of polarised, emotionally charged interactions—even in the presence of exposure to opposing views—suggests that efforts to foster consensus around climate policy must account for the affective and identity-driven nature of public discourse. Moreover, the asymmetric structure of negativity and the association between political alignment and information reliability can contribute to public confusion, distrust in science, and a fragmentation of climate narratives, making it harder to establish the shared understanding necessary for effective action. Addressing polarisation in this context requires more than correcting low-reliability information—it demands strategies that bridge ideological divides and rebuild trust in both institutions and science.
Our static analysis of the polarisation around climate change could be extended to assess how it evolves with time as a function of internal dynamics, external sources, and the unfolding of events. For instance, coordinated campaigns have been observed to impact social media dynamics [84], and could play a role in shaping the emergence or intensification of polarisation. It is unclear whether message negativity is intrinsic to the topic or evolves with the discussion, potentially triggered by specific interactions. The asymmetric spread of emotionally polarised content suggests that by targeting specific small groups, polarisation could be significantly decreased [80]. However, the opacity of most online social networks makes content moderation a major challenge. It would be essential to assess whether content moderation or fact-checking could have a positive impact on the reduction of polarisation. Since our results suggest that individuals sharing low-reliability information are exposed to opposing views, the development of intervention strategies becomes a major challenge. The role of exposure in increasing polarisation could be further addressed by investigating the temporal evolution of user ideology and reliability as a function of their progressive interactions with others. In a similar fashion, it opens the door to the development of models to unveil how the exposure to opposing views can further exacerbate polarisation.
Data availability statement
The Twitter data are made available in accordance with Twitter’s terms of service at https://github.com/clint-project/clint-data (https//doi.org:/10.5281/zenodo.14095218) and the codes to calculate the metrics at https://github.com/clint-project/polarization_metrics (https://doi.org/10.5281/zenodo.14095108). The YouTube data is available at
https://doi.org:/10.5281/zenodo.15046657.
Author contribution
A.B: data collection, data curation, investigation, methodology, visualization; J.M.: data collection, data curation, investigation, methodology. E.C.: conceptualization, investigation, methodology, writing; J.V.: conceptualization, investigation, methodology, writing. All authors gave final approval for publication and agreed to be held accountable for the work performed therein.
Acknowledgements
J.M. was a fellow of Eurecat’s “Vicente López” PhD grant program. This work was financially supported by the Catalan Government through the funding grant ACCIÓ-Eurecat (Project TRAÇA 2023 - CLINT). E.C. acknowledges support from the Spanish grants PGC2018-094754-B-C22 and PID2021-128005NB-C22, funded by MCIN/AEL 10.13039/501100011033 and ”ERDF A way of making Europe”, and from Generalitat de Catalunya under project 2021-SGR-00856.
Appendix A Dataset collection and description
In this paper, we employed Twarc, a tool for collecting data from Twitter, to delve into user-generated content. By utilising hashtags as search queries, particularly the search and stream API, we collected Twitter conversations, targeting specific themes and topics relevant to our study (COP27, IPCC, Climate Refugees, and Doñana). This methodological approach facilitated the systematic collection of rich and diverse data. These are the specific hashtags that we utilised to build the datasets:
  • COP27: COP27, TogetherForImplementation, COPTV, PRECOP, PRECOP27, LossAndDamage, Egypt_COP27, COP27Egypt, ClimateAction, ClimateCrisis, ClimateJustice.
  • IPCC: ipcc, @ipcc_ch.
  • Climate Refugees: climate (refugees OR refugee), climate migration, climate displaced.
  • Doñana: Doñana.
We also provide the detailed json queries used in the extraction in Listings LABEL:lst:cop27, LABEL:lst:ipcc, LABEL:lst:climate-refugees and LABEL:lst:donana.
Listing 1: COP27 Query Parameters
1{
2 "start-time": "2022-09-01",
3 "end-time": "2022-11-28",
4 "keywords": [
5 "COP27",
6 "TogetherForImplementation",
7 "COPTV",
8 "PRECOP",
9 "PRECOP27",
10 "LossAndDamage",
11 "Egypt_COP27",
12 "COP27Egypt",
13 "ClimateAction",
14 "ClimateCrisis",
15 "ClimateJustice"
16 ]
17}
Listing 2: IPCC Query Parameters
1{
2 "start-time": "2023-03-18",
3 "end-time": "2023-03-27",
4 "keywords": [
5 "ipcc",
6 "@ipcc_ch"
7 ]
8}
Listing 3: Climate Refugees Query Parameters
1[
2 {
3 "start-time": "2008-01-01",
4 "end-time": "2023-01-01",
5 "keywords": [
6 "climate (refugees OR refugee)",
7 "climate migration",
8 "climate displaced"
9 ]
10 }
11]
Listing 4: Doñana Query Parameters
1{
2 "start-time": "2019-01-01",
3 "end-time": "2023-05-01",
4 "keywords": [
5 "Do~nana"
6 ]
7}
The YouTube videos appearing in the Twitter dataset were filtered using specific keywords for each dataset. The keywords used were as follows: COP27 – ”COP27”, ”climate”, and ”COP”; IPCC – ”IPCC”, ”report”, and ”climate”; Climate Refugees – ”Refugee” and ”climate”; and Doñana – ”Doñana” and ”clima.”
In Fig. S1 and Fig. S2 we show the timeseries of each dataset for Twitter and YouTube, respectively.
Figure S1: Timeseries of the COP27, IPCC, Climate Refugees and Doñana datasets. Daily timeseries of tweets in the datasets gathered for COP27, IPCC, Climate Refugees and Doñana.
Figure S2: Timeseries of comments in the Youtube datasets. Number of comments per date in the Youtube datasets of COP27, IPCC, Climate Refugees and Doñana. For visualisation purposes, we show the comments after 2021.
Appendix B Network analysis and structural polarisation
S2 Influencers by community
In this section, we identify the top influencers within different communities for each of the key datasets: COP27, IPCC, Climate Refugees, and Doñana. The influencers are ranked by their total inflow of interactions, highlighting those users who have the most significant impact within their respective communities. Tables S1, S2, S3, and S4 present the top 4 most relevant users in the ten largest communities for each dataset.
User Community Pagerank Inflow In-degree Outflow Out-degree Flow ratio Degree ratio
MikeHudema 0 0.004604 126323 42099 12 8 0.999905 0.999810
COP27P 0 0.008392 108582 45664 380 142 0.996513 0.996900
antonioguterres 0 0.006937 79507 38380 5 5 0.999937 0.999870
UN 0 0.003134 43045 24757 211 56 0.995122 0.997743
JamesMelville 1 0.007062 49172 31076 8 5 0.999837 0.999839
DiEM_25 1 0.005087 28848 28496 11 6 0.999619 0.999789
BernieSpofforth 1 0.002298 26501 15218 21 10 0.999208 0.999343
DrEliDavid 1 0.000952 13377 11137 0 0 1.000000 1.000000
kimpaim 2 0.001172 24266 18407 2 2 0.999918 0.999891
TerraBrasilnot 2 0.001122 16518 12962 0 0 1.000000 1.000000
VemPraRua_br 2 0.000899 16234 13991 0 0 1.000000 1.000000
marciolabre 2 0.000652 13147 13028 0 0 1.000000 1.000000
LulaOficial 3 0.010655 130064 38689 24 10 0.999816 0.999742
choquei 3 0.005587 25475 14922 1 1 0.999961 0.999933
Reuters 3 0.001247 12269 8240 67 13 0.994569 0.998425
MarinaSilva 3 0.002625 11904 8575 11 6 0.999077 0.999301
CarolineLucas 4 0.001137 33989 14931 40 31 0.998825 0.997928
NicolaSturgeon 4 0.000565 10778 4993 59 25 0.994556 0.995018
_david_ho_ 4 0.000498 6706 6246 8 8 0.998808 0.998721
johnestevens 4 0.000809 6553 6154 2 2 0.999695 0.999675
RishiSunak 5 0.005127 37770 24640 3 3 0.999921 0.999878
EmmanuelMacron 5 0.003232 16585 12178 2 2 0.999879 0.999836
Poulin2012 5 0.001389 12919 12735 3 3 0.999768 0.999764
MickaCorreia 5 0.000466 3869 3054 45 19 0.988503 0.993817
POTUS 6 0.004068 24808 15824 1 1 0.999960 0.999937
SpeakerPelosi 6 0.000971 $94$ 5361 9 6 0.999053 0.998882
JustinTrudeau 6 0.000316 5750 4247 2 2 0.999652 0.999529
washingtonpost 6 0.001279 5564 4655 14 2 0.997490 0.999571
Novozymes 7 0.000194 7735 3629 16 10 0.997936 0.997252
AfDB_Group 7 0.000376 5991 2294 358 104 0.943613 0.956631
RockefellerFdn 7 0.000438 5845 3381 346 118 0.944112 0.966276
WilliamsRuto 7 0.000691 4941 3163 9 4 0.998182 0.998737
Danielbricen 8 0.001689 12626 7694 4 2 0.999683 0.999740
IvanDuque 8 0.000791 9105 6317 1 1 0.999890 0.999842
alertaLatam 8 0.000337 6695 4717 17 3 0.997467 0.999364
marcelamvzyya 8 0.000140 5918 3507 0 0 1.000000 1.000000
petrogustavo 9 0.005749 54899 24016 23 14 0.999581 0.999417
infopresidencia 9 0.001116 18209 6575 60 17 0.996716 0.997421
FranciaMarquezM 9 0.000399 10795 5306 22 17 0.997966 0.996806
GustavoBolivar 9 0.001154 8534 6545 7 2 0.999180 0.999695
Table S1: Top 4 most relevant users by total inflow in the ten largest communities of the COP27 dataset
User Community Pagerank Inflow In-degree Outflow Out-degree Flow ratio Degree ratio
IPCC_CH 0 0.173282 86702 48491 22 12 0.999746 0.999753
ed_hawkins 0 0.002570 3194 2119 60 29 0.981561 0.986499
MrMatthewTodd 0 0.001954 2766 2474 54 37 0.980851 0.985265
CharlieJGardner 0 0.002120 2720 2297 5 4 0.998165 0.998262
uksciencechief 1 0.004402 12700 5452 1 1 0.999921 0.999817
BernieSpofforth 1 0.005600 5735 4972 0 0 1.000000 1.000000
fmeeus1 1 0.004689 3973 1800 15 8 0.996239 0.995575
Lauratobin1 1 0.000858 3067 1287 4 3 0.998697 0.997674
jasonhickel 2 0.008178 4249 4120 2 2 0.999530 0.999515
greenpeace_esp 2 0.003291 2092 1986 28 11 0.986792 0.994492
millares 2 0.001390 1788 1786 5 4 0.997211 0.997765
ionebelarra 2 0.000286 592 444 3 3 0.994958 0.993289
GretaThunberg 3 0.007815 8539 8115 0 0 1.000000 1.000000
CarolineLucas 3 0.003437 5833 3971 5 5 0.999144 0.998742
paulpowlesland 3 0.001666 1189 1179 0 0 1.000000 1.000000
guardian 3 0.000646 1016 915 0 0 1.000000 1.000000
KHayhoe 4 0.006960 9241 7626 82 39 0.991205 0.994912
SenSchumer 4 0.001571 4078 1973 2 2 0.999510 0.998987
curious_founder 4 0.004436 2067 1779 9 5 0.995665 0.997197
remblance_erin 4 0.001244 1380 1325 9 9 0.993521 0.993253
rahmstorf 5 0.003506 3135 2032 35 17 0.988959 0.991703
micha_bloss 5 0.007133 2687 2276 3 2 0.998885 0.999122
Martin_Bethke 5 0.001183 2124 1222 71 40 0.967654 0.968304
HolzheuStefan 5 0.001040 1787 1241 26 13 0.985659 0.989633
antonioguterres 6 0.015073 6512 4285 1 1 0.999846 0.999767
UN 6 0.007672 4093 2753 8 2 0.998049 0.999274
BarackObama 6 0.002001 1794 1775 0 0 1.000000 1.000000
HillaryClinton 6 0.001469 1398 1375 0 0 1.000000 1.000000
JKSteinberger 7 0.006622 2121 1930 76 26 0.965407 0.986708
frantecol 7 0.001867 2101 2052 6 4 0.997152 0.998054
jennystojkovic 7 0.000888 1519 1324 8 4 0.994761 0.996988
NLRebellion 7 0.000375 571 407 26 17 0.956449 0.959906
MikeHudema 8 0.003279 5639 3757 1 1 0.999823 0.999734
s_guilbeault 8 0.001042 2369 1082 4 2 0.998314 0.998155
tveitdal 8 0.000368 561 409 35 7 0.941275 0.983173
gmbutts 8 0.000383 512 406 4 3 0.992248 0.992665
DrLuetke 9 0.000737 741 719 0 0 1.000000 1.000000
Axel_Bojanowski 9 0.001738 561 406 11 3 0.980769 0.992665
welt 9 0.000674 543 471 2 1 0.996330 0.997881
Schlautropf 9 0.000378 419 375 14 12 0.967667 0.968992
Table S2: Top 4 most relevant users by total inflow in the ten largest communities of the IPCC dataset
User Community Pagerank Inflow In-degree Outflow Out-degree Flow ratio Degree ratio
AOC 0 0.012350 35741 31238 9 7 0.999748 0.999776
BernieSanders 0 0.002378 12756 11197 2 2 0.999843 0.999821
propublica 0 0.003287 12325 8952 87 33 0.992991 0.996327
POTUS 0 0.003139 8389 7316 0 0 1.000000 1.000000
MikeHudema 1 0.002022 12798 10474 13 8 0.998985 0.999237
AssaadRazzouk 1 0.002807 10615 7388 44 31 0.995872 0.995822
PaulEDawson 1 0.000761 10559 5675 15 8 0.998581 0.998592
ClimateBen 1 0.003454 10327 9121 135 62 0.987096 0.993248
zoenone0none 2 0.011598 24876 24855 3 2 0.999879 0.999920
studentactivism 2 0.011936 19994 19861 2 2 0.999900 0.999899
ajplus 2 0.009282 6795 5790 18 6 0.997358 0.998965
femalekissinger 2 0.001781 5197 5195 1 1 0.999808 0.999808
Refugees 3 0.009205 29470 15697 197 86 0.993360 0.994551
UN 3 0.006119 16457 12513 91 33 0.994501 0.997370
UNmigration 3 0.001623 8475 4151 278 113 0.968239 0.973499
antonioguterres 3 0.002124 6676 5626 0 0 1.000000 1.000000
nytimes 4 0.009325 21229 18971 51 16 0.997603 0.999157
guardian 4 0.002638 7924 6603 9 4 0.998865 0.999395
LeoDiCaprio 4 0.001235 5447 4897 9 9 0.998350 0.998166
NatGeo 4 0.001027 5283 4718 0 0 1.000000 1.000000
MrAhmednurAli 5 0.002175 11799 11703 1 1 0.999915 0.999915
jeremycorbyn 5 0.003088 6753 5402 3 2 0.999556 0.999630
LaylaMoran 5 0.000665 3725 3711 0 0 1.000000 1.000000
bmay 5 0.000566 3383 3029 15 13 0.995586 0.995726
BreitbartNews 6 0.000559 3076 2481 0 0 1.000000 1.000000
MaximeBernier 6 0.000966 2799 2022 17 11 0.993963 0.994589
PrisonPlanet 6 0.001274 1912 1711 1 1 0.999477 0.999416
Liz_Wheeler 6 0.000363 1810 1787 0 0 1.000000 1.000000
Kon__K 7 0.001581 9741 6362 46 24 0.995300 0.996242
bruce_haigh 7 0.000417 4512 2721 115 86 0.975146 0.969362
simonahac 7 0.000289 2775 2060 58 47 0.979527 0.977693
abcnews 7 0.000239 2051 1671 2 2 0.999026 0.998805
JustinTrudeau 8 0.003543 15733 15066 5 5 0.999682 0.999668
ianbremmer 8 0.000955 4529 4469 9 4 0.998017 0.999106
cjwerleman 8 0.000434 1052 1017 3 3 0.997156 0.997059
MusaNV18 8 0.000294 931 877 4 3 0.995722 0.996591
ndcodeine 9 0.006104 24228 24205 0 0 1.000000 1.000000
RoadtoMUT2022 9 0.000038 157 157 0 0 1.000000 1.000000
chloeprw 9 0.000016 71 71 0 0 1.000000 1.000000
themomentumco 9 0.000009 40 40 0 0 1.000000 1.000000
Table S3: Top 4 most relevant users by total inflow in the ten largest communities of the Climate Refugees dataset
User Community Pagerank Inflow In-degree Outflow Out-degree Flow ratio Degree ratio
JuanMa_Moreno 0 0.014113 37135 14986 25 8 0.999327 0.999466
alfonso_ussia 0 0.003464 17027 9717 41 30 0.997598 0.996922
okdiario 0 0.003296 13310 5792 44 11 0.996705 0.998104
AndaluciaJunta 0 0.011000 11794 6983 386 66 0.968309 0.990637
WWFespana 1 0.006797 20837 8588 1292 201 0.941615 0.977131
juralde 1 0.005168 19984 9507 363 94 0.982160 0.990209
Santi_MBarajas 1 0.003885 19234 7772 270 65 0.986157 0.991706
SEO_BirdLife 1 0.004740 10059 4557 499 138 0.952737 0.970607
FonsiLoaiza 2 0.009273 13695 9416 5 1 0.999635 0.999894
JA_DelgadoRamos 2 0.002128 10446 5413 23 12 0.997803 0.997788
Toni_Valero 2 0.002191 9947 4700 64 27 0.993607 0.994288
JavierArocaA 2 0.001303 8632 4994 84 30 0.990363 0.994029
TeresaRodr_ 3 0.002844 10018 5616 209 73 0.979564 0.987168
Ainhoasauria 3 0.003481 5677 5676 0 0 1.000000 1.000000
MiguelMorenatti 3 0.001572 4024 3916 5 4 0.998759 0.998980
lavozdelsures 3 0.000936 3343 2414 475 191 0.875589 0.926679
carlosromeroco 4 0.016908 51995 16142 4210 302 0.925096 0.981635
PartidoPACMA 4 0.004616 10101 8238 32 15 0.996842 0.998182
fromerofoto 4 0.001280 4799 1132 887 90 0.844003 0.926350
guardiacivil 4 0.004166 3612 2798 7 2 0.998066 0.999286
sanchezcastejon 5 0.032978 67372 26401 2 2 0.999970 0.999924
eldiarioes 5 0.008312 24556 11979 205 32 0.991721 0.997336
PSOE 5 0.003014 22664 7498 47 15 0.997931 0.998003
iescolar 5 0.006451 21280 10952 6 3 0.999718 0.999726
jcanadellb 6 0.000574 1478 1084 15 8 0.989953 0.992674
sninobecerra 6 0.000243 789 739 0 0 1.000000 1.000000
puntocriticoDH 6 0.000112 675 201 358 16 0.653437 0.926267
jm_clavero 6 0.000113 528 439 1 1 0.998110 0.997727
ClimateBen 7 0.002295 798 740 2 1 0.997500 0.998650
WWF_Deutschland 7 0.000356 246 199 14 5 0.946154 0.975490
iacbe 7 0.000101 173 108 13 7 0.930108 0.939130
AlanDaviesbirds 7 0.000195 158 107 22 10 0.877778 0.914530
ParlamentoAnd 8 0.002065 2855 2335 24 10 0.991664 0.995736
CiudadanosCs 8 0.000200 724 501 24 14 0.967914 0.972816
Cs_Andalucia 8 0.000254 557 440 67 32 0.892628 0.932203
AsocParqueDunar 8 0.000136 430 114 101 26 0.809793 0.814286
GotTalentES 9 0.001144 1618 1617 0 0 1.000000 1.000000
proalmerienses 9 0.000008 22 21 0 0 1.000000 1.000000
RaholaOficial 9 0.000007 22 13 16 16 0.578947 0.448276
chiquisanz 9 0.000001 10 5 308 60 0.031447 0.076923
Table S4: Top 4 most relevant users by total inflow in the ten largest communities of the Doñana dataset
S2 Language analysis
We present the most frequently used languages across the largest communities for each dataset in Tables S5, S6, S7, and S8, which correspond to the COP27, IPCC, Climate Refugees, and Doñana datasets, respectively.
Community Top 1 language Top 1 language ratio Top 2 language Top 2 language ratio
0 en 0.911 de 0.051
1 en 0.860 de 0.060
2 pt 0.980 es 0.011
3 pt 0.927 en 0.047
4 en 0.994 fr 0.001
5 fr 0.871 en 0.113
6 en 0.917 pl 0.021
7 en 0.911 fr 0.026
8 es 0.945 en 0.042
9 es 0.926 en 0.061
Table S5: Most used languages in the ten largest communities of the COP27 dataset
Community Top 1 language Top 1 language ratio Top 2 language Top 2 language ratio
0 en 0.985 de 0.003
1 en 0.773 nl 0.180
2 es 0.699 en 0.261
3 en 0.993 de 0.005
4 en 0.985 de 0.005
5 de 0.930 en 0.064
6 en 0.986 es 0.004
7 en 0.647 nl 0.323
8 en 0.791 sv 0.192
9 de 0.955 en 0.043
Table S6: Most used languages in the ten largest communities of the IPCC dataset
Community Top 1 language Top 1 language ratio Top 2 language Top 2 language ratio
0 en 0.998 ro 0.001
1 en 0.987 fi 0.003
2 en 0.990 id 0.009
3 en 0.990 es 0.003
4 en 0.991 it 0.004
5 en 0.998 it 0.001
6 en 0.990 nl 0.004
7 en 0.997 fr 0.002
8 en 0.996 it 0.001
9 th 0.991 en 0.009
Table S7: Most used languages in the ten largest communities of the Climate Refugees dataset
Community Top 1 language Top 1 language ratio Top 2 language Top 2 language ratio
0 es 0.996 pt 0.001
1 es 0.907 en 0.077
2 es 0.993 ca 0.003
3 es 0.969 pt 0.015
4 es 0.952 pt 0.024
5 es 0.991 en 0.005
6 es 0.635 ca 0.295
7 en 0.609 de 0.283
8 es 0.978 en 0.011
9 es 0.998 pt 0.001
Table S8: Most used languages in the ten largest communities of the Doñana dataset
The language distribution across the largest communities in the COP27, IPCC, Climate Refugees, and Doñana datasets reveals notable patterns. These patterns not only reflect the diversity of linguistic representation within each dataset but also provide insights into the global reach.
S2 Analysis of global structural polarisation
To assess the degree of polarisation within the online discourse surrounding key climate-related topics, we calculate the standardised polarisation in the COP27, IPCC, Climate Refugees, and Doñana networks, following the method outlined in [34]. The polarisation is standardised by comparing the observed polarisation $\Phi(G)$ against the expected polarisation $\Phi(G_{CM})$ from a corresponding null model, resulting in a z-score that accounts for network-specific variations. The formula for this calculation is given by:
$\hat{\Phi}_{z}(G)=\frac{\Phi(G)-\Phi(G_{CM})}{\sqrt{\langle\Phi(G_{CM})^{2}\rangle-\langle\Phi(G_{CM})\rangle^{2}}}.$ (S1)
Figure S3 presents the standardised polarisation across the retweet networks for the mentioned datasets, illustrating the extent of polarisation after denoising and z-score calculation.
Figure S3: Standardised polarisation in the Twitter dataset. Standardized measures of polarisation in the retweet networks after the denoising and the calculation of z-scores.
Furthermore, we have computed in Fig. S4 the polarisation scores in the subgraph between the two largest communities with opposing bias that speak the same language.
Figure S4: Standarized polarisation in Twitter by interaction type. Standardized modularity and the E-I indices in (a) the retweet and (b) mention networks. We computed the polarisation over the subnetwork of the two largest communities of opposed political bias. We have calculated the indices in the real network, in 100 realizations of the configuration model with preserved degree sequences, and standardized them.
S2 Bias and reliability
Figure S5: Distribution of bias and reliability in the largest communities.(a-d) Distribution of bias and reliability in the seven largest communities for (a) COP27, (b) IPCC, (c) Climate Refugees and (d) Doñana.
We tested the significance of the observed bias (Fig. S6) and reliability (Fig. S7) to determine whether the differences identified in our analysis are statistically meaningful. In both cases, the differences between the two selected communities in each network are statistically significant.
Figure S6: Mann-Whitney U test between the bias of the top 7 communities. Result of the Mann-Whitney U test between the bias of the top 7 communities. The test was conducted to assess if the community in the vertical axis is larger than the community in the horizontal axes. Asterisks report the significance of the tests (*p-value<<0.05, **p-value<<0.01 and ***p-value<<0.001).
Figure S7: Mann-Whitney U test between the reliability of the top 7 communities. Result of the Mann-Whitney U test between the reliability of the top 7 communities. The test was conducted to assess if the community in the vertical axis is larger than the community in the horizontal axes. Asterisks report the significance of the tests (*p-value<<0.05, **p-value<<0.01 and ***p-value<<0.001).
S2 Isolation and entropy
To evaluate the isolation and interaction of communities, we calculated the directed interactions between them, and we computed the normalised outflow difference given by
$\Delta F^{out}_{i}=\frac{\sum_{j,j\neq i}T_{ij}-T_{ii}}{\sum_{\forall j}T_{ij}},$ (S2)
and the entropy of flows as
$H^{out}_{i}=-\sum_{\forall j}p_{ij}logp_{ij},$ (S3)
where $T_{ij}$ is the number of interactions from community $i$ to community $j$ and $p_{ij}$ is the probability of having a retweet from $i$ to $j$ given by $T_{ij}/\sum_{\forall j}T_{ij}$. While the first index quantifies the ratio between external and internal links and goes from 1-1 to $1$ when all flows are external, the entropy provides information on the variety of those flows. The entropy is 0 when there is no variety of flows and $1$ when it is maximum. In Fig. S8, we represent the outflow entropy as a function of the isolation for the four networks.
We computed the minimum and maximum attainable entropy values as a function of $\Delta F^{out}_{i}$ and $p_{ii}$ for orientation. Given a value of $\Delta F^{out}_{i}$ we have
$\frac{\sum_{j,j\neq i}T_{ij}}{{\sum_{\forall j}}T_{ji}}=\Delta F_{i}+\frac{T_{ii}}{{\sum_{\forall j}T_{ij}}}=\Delta F_{i}+p_{ii}.$ (S4)
The minimum value of the entropy corresponds to the situation when a community is connected only to itself and one external community (except for the extreme case $\Delta F_{i}=1$), which we can plug into the entropy as
$H^{in}_{i}=-p_{ii}logp_{ii}-\Delta F_{i}+p_{ii}log\Delta F_{i}+p_{ii}.$ (S5)
The maximum value of the entropy appears when a community is linked to all the others with a flow $T_{ij}/(N_{c}-1)$, where $N_{c}$ is the total number of communities in the network. Thus, it can be written as
$H^{in}_{i}=-p_{ii}logp_{ii}-\sum^{N_{c}-1}_{j=0}\frac{\Delta F_{i}+p_{ii}}{N_{c}-1}log\frac{\Delta F_{i}+p_{ii}}{N_{c}-1}.$ (S6)
In Figs. S8-S9 we show the outflow entropy as a function of the outgoing isolation for the retweet and mention networks between communities.
Figure S8: Community connectivity and entropy in retweet networks. Outflow entropy as a function of the normalised outgoing isolation in the retweet networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of feasible values.
Figure S9: Community connectivity and entropy in mention networks. Outflow entropy as a function of the normalised outgoing connectivity in the mention networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of attainable values.
We also show in Fig. S10 the inflow entropy as a function of the incoming isolation for the retweet network between communities.
Figure S10: Community connectivity and entropy in retweet networks. Inflow entropy as a function of the normalized incoming connectivity in the retweet networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of attainable values.
We report in Tables S9 and S10 we report the values of isolation and entropy for the 8 largest communities in each network.
Network Interaction Community $\Delta F^{out}$ $H^{out}$ Interaction Community $\Delta F^{out}$ $H^{out}$
0 -0.68 0.14 0 -0.57 0.17
1 -0.98 0.01 1 0.45 0.24
2 -0.72 0.10 2 0.03 0.15
IPCC Retweet 3 -0.61 0.12 Mention 3 0.73 0.19
4 -0.50 0.16 4 0.83 0.12
5 -0.76 0.09 5 0.26 0.24
6 -0.60 0.14 6 0.58 0.15
7 -0.43 0.18 7 0.51 0.18
0 -0.76 0.10 0 -0.50 0.17
1 -0.88 0.05 1 0.37 0.25
2 -0.97 0.01 2 0.18 0.19
COP27 Retweet 3 -0.94 0.03 Mention 3 -0.78 0.08
4 -0.60 0.10 4 0.09 0.22
5 -0.75 0.09 5 -0.32 0.18
6 -0.50 0.14 6 0.04 0.21
7 -0.51 0.13 7 0.01 0.20
0 -0.95 0.03 0 0.28 0.20
1 -0.48 0.16 1 -0.02 0.25
2 -0.33 0.17 2 0.55 0.27
Doñana Retweet 3 -0.33 0.18 Mention 3 0.47 0.29
4 -0.59 0.13 4 -0.31 0.19
5 -0.49 0.14 5 0.08 0.23
6 -0.32 0.19 6 0.70 0.24
7 -0.82 0.06 7 0.20 0.23
0 -0.67 0.10 0 -0.05 0.24
1 -0.58 0.13 1 0.24 0.27
2 -0.71 0.08 2 0.55 0.30
Climate Refugees Retweet 3 -0.68 0.10 Mention 3 -0.50 0.16
4 -0.57 0.12 4 0.33 0.26
5 -0.62 0.11 5 0.07 0.25
6 -0.93 0.03 6 -0.13 0.21
7 -0.72 0.08 7 -0.55 0.13
Table S9: Isolation and entropy metrics for the IPCC, COP27, Doñana, and Climate Refugees datasets, categorised by retweet and mention interactions for each community in terms of outflows.
Network Interaction Community $\Delta F^{in}$ $H^{in}$ Interaction Community $\Delta F^{in}$ $H^{in}$
0 -0.60 0.17 0 0.29 0.39
1 -0.97 0.02 1 -0.68 0.12
2 -0.73 0.12 2 -0.73 0.12
IPCC Retweet 3 -0.48 0.18 Mention 3 0.72 0.32
4 -0.35 0.21 4 0.60 0.33
5 -0.91 0.05 5 0.16 0.22
6 -0.63 0.14 6 0.74 0.35
7 -0.55 0.16 7 0.17 0.23
0 -0.69 0.12 0 -0.28 0.25
1 -0.83 0.07 1 0.06 0.22
2 -0.98 0.01 2 -0.68 0.08
COP27 Retweet 3 -0.93 0.03 Mention 3 -0.58 0.13
4 -0.69 0.09 4 -0.02 0.23
5 -0.82 0.07 5 0.25 0.29
6 -0.46 0.16 6 0.49 0.33
7 -0.54 0.14 7 0.11 0.25
0 -0.95 0.03 0 0.28 0.26
1 -0.41 0.17 1 -0.18 0.23
2 -0.48 0.15 2 0.02 0.24
Doñana Retweet 3 -0.36 0.18 Mention 3 0.40 0.29
4 -0.60 0.14 4 -0.31 0.21
5 -0.38 0.16 5 0.38 0.24
6 -0.49 0.15 6 0.41 0.28
7 -0.86 0.05 7 0.16 0.23
0 -0.65 0.11 0 0.15 0.27
1 -0.49 0.15 1 -0.02 0.26
2 -0.70 0.10 2 0.40 0.27
Climate Refugees Retweet 3 -0.67 0.11 Mention 3 -0.24 0.22
4 -0.52 0.14 4 0.70 0.32
5 -0.70 0.10 5 -0.10 0.23
6 -0.96 0.02 6 -0.21 0.18
7 -0.87 0.05 7 -0.71 0.09
Table S10: Isolation and entropy metrics for the IPCC, COP27, Doñana, and Climate Refugees datasets, categorised by retweet and mention interactions for each community in terms of inflows.
S2 Echo chambers
Echo chambers emerge in online discourse when users predominantly interact with like-minded individuals, reinforcing their existing beliefs. In political discussions, this phenomenon can manifest in different ways depending on the type of interaction network. In the mention network, users tend to engage more with others who share similar political biases, forming bias-based echo chambers. Meanwhile, in the retweet network, clustering occurs based on the reliability of the media sources that users amplify, leading to reliability-based echo chambers. These dynamics are analysed across key climate-related topics, including COP27, IPCC, Climate Refugees, and Doñana. Figures S11 and S12 illustrate these patterns, highlighting the structure and polarisation of online conversations.
Figure S11: Detection of echo chambers in the mention network for political bias. Political bias of the neighbourhood as a function of the bias of a user in the mention network for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
Figure S12: Detection of echo chambers in the retweet network for reliability. Media reliability of the neighbourhood as a function of the bias of a user in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. studied by considering only retweets.
Figure S13: Chamber overlap distribution $q_{ij}$. Chamber overlap distribution for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana calculated on the top 2020 users of each community.
The chamber overlap can distinguish between users sharing similar echo chambers (Fig.S14). We constructed the user-to-user matrices where the values correspond to their chamber overlap. Using a hierarchical agglomerative algorithm, we clustered the distance matrix and obtained an almost perfect separation of users by communities. The separation between the users in the communities is neater in the COP27 and Doñana datasets compared to IPCC and Climate Refugees. Nevertheless, we observe a clear trend of similar echo chambers among users within the same community. In the COP27 network, the pattern is more evident for the users in the mainstream community (in blue). The separation between communities and the matching of echo chambers suggest a clearer separation in the Doñana dataset.
Figure S14: Chamber overlap matrix between users clustered using a hierarchical agglomerative algorithm. Chamber overlap between the $20$ users with more retweets in the two largest communities with opposed bias in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
Appendix C Policy and affective polarisation
S3 Hashtag analysis
Understanding the dynamics of hashtag usage across different communities provides valuable insights into the discourse surrounding key topics. In Fig. S15, we present the 20 most tweeted and retweeted hashtags for the two communities analyzed in each network. Our findings reveal a substantial overlap in hashtag usage between communities, with a dominance of generalistic hashtags. To further investigate these patterns, we assess the significance of differences in hashtag presence in Fig. S16. The results indicate that, in all cases, the variation in hashtag abundance between communities is statistically significant, highlighting distinct engagement patterns despite the observed overlap.
Figure S15: Ranking of hashtags by number of interactions in the top two opposing communities. Ranking of hashtags in the two largest communities with opposed political bias in the (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks.
Figure S16: Hashtag proportion test. We have conducted the ztest between the proportion of hashtags in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks. Blue bars indicate significance (p-value<<0.05).
S3 Topic and sentiment analysis
We have implemented a Latend Dirichlet Allocation model (LDA) to uncover the main topics discussed in each of the datasets. In Fig. S17 we report the most important words for each of the topics detected in the datasets. We have analyzed the polarisation around each of the topics Fig. S18.
Figure S17: Topics detected in the networks analized. Results of the Latent Dirichlet Allocation model (LDA) for eight topics in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks.
Figure S18: Distribution of the polarisation around the discussion topics. Topics detected and polarisation in (a) COP27, (b) IPCC,(c) Climate Refugees, (d) Doñana. Blue bars indicate that the frequency of topics between communities is statistically significant.
We have implemented a sentiment analysis on the four Twitter datasets using a Roberta-based model fine-tuned on around 58 million of Tweets [77]. We provide the Mann-Whitney U tests on the quoted messages in Fig. S19.
Figure S19: Mann-Whitney U test in quote sentiment. We have conducted the Mann-Whitney U test of the negative sentiment between communities in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks. The test was conducted to assess if the entries in the vertical axes are greater than those in the horizontal axes.
Appendix D Cross-platform analysis
We present the network of YouTube posts extracted from the Twitter datasets, where each node represents a post and is linked based on the number of shared commenters Figure S20. Nodes are color-coded according to community structures identified through modularity optimization [69]. To further investigate the network’s structure, we analyze a subnetwork focusing on communities that predominantly spread skeptical narratives about climate change, as well as the largest community associated with mainstream content. The results of this analysis are depicted in Fig. S21.
Figure S20: Network of Youtube posts extracted from the Twitter datasets. Each node corresponds to a post and links to the number of co-commenting users. Each node is coloured according to the community calculated by optimising the modularity [69].
Figure S21: Network of YouTube posts after selecting the two main communities of mainstream and low-reliability content. Interaction between the posts after selecting the larger communities with mainstream and low-reliability content. We calculated the communities by optimizing the modularity for two partitions. We identified the positioning of the communities by inspecting the videos.
In Fig S22 we show the Twitter subnetwork of users that shared YouTube links related to climate change coloured according to the YouTube community of the videos. Only connected components with more than 5 users are depicted.
Figure S22: Network between Twitter users coloured by YouTube community. Interactions between Twitter users that have published YouTube Content in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. Each dot is coloured by the content of YouTube communities they have tweeted on.
We have computed the probability $P(C_{i},C_{j})$ of finding a user that posted a video in YouTube community $C_{i}$ connected to a user that posted a video in YouTube community $C_{j}$. To account for network effects, we have calculated $\widetilde{P}(C_{i},C_{j})=P(C_{i},C_{j})/P^{\rm null}(C_{i},C_{j})$ normalizing the observed probabilities by the corresponding null values over 100 random reshufflings of YouTube communities (Fig. S23)
Figure S23: YouTube Community correlations in the Twitter network. Probability of finding a user than posted a video in a YouTube community $C_{i}$ next to a user that posted a video in a YouTube community $C_{j}$. The results are normalized by 1000 random reshufflings of YouTube communities. The results correspond to (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana networks.
network community channel name viewCount subscriberCount videoCount
cop27 0 ABC News 14083290454 15900000 86239
cop27 0 TEDx Talks 7712122864 40100000 210266
cop27 0 WION 4701543170 8470000 136298
cop27 0 Al Jazeera English 4158638169 12800000 111285
cop27 1 ABC News 14083290454 15900000 86239
cop27 1 BBC News Hindi 7893209075 17800000 24225
cop27 1 BBC News 5099185254 15600000 20649
cop27 1 Hindustan Times 4842818648 6530000 58180
cop27 2 WION 4701543170 8470000 136298
cop27 2 Sky News Australia 4140464784 3900000 137500
cop27 2 CNN-News18 2967286345 6660000 160062
cop27 2 euronews 1368026874 2320000 99119
cop27 3 Democracy Now! 672210420 1980000 15630
cop27 3 BreakThrough News 138428175 609000 1403
cop27 3 CNBC International TV 116672225 362000 23849
cop27 3 Earthling Ed 54396530 452000 333
cop27 4 UOL 3147321219 3880000 62664
cop27 4 CNN Brasil 2660548853 4780000 79898
cop27 4 Poder360 441840465 1300000 19396
cop27 4 United Nations 406648137 2870000 16539
cop27 5 DW Español 2280313778 4150000 39964
cop27 5 EL PAÍS 1525037481 2550000 32904
cop27 5 RTVE Noticias 1378481206 2020000 25701
cop27 5 Presidencia de la República - Colombia 92079186 318000 20699
donana 0 Negocios TV 961486669 1010000 51192
donana 0 COPE 428257253 680000 6994
donana 0 LibertadDigital 332062887 684000 14666
donana 0 DISTRITOTV 172849322 331000 17512
donana 1 Canal Red 40410031 309000 700
donana 1 Pandemia Digital 21969182 77300 1701
donana 2 EL PAÍS 1525037481 2550000 32904
donana 2 RTVE Noticias 1378481206 2020000 25701
donana 2 elDiarioes 172800438 320000 9763
donana 2 Hoy por Hoy 37690183 71900 2893
ipcc 0 ABC News 14083290454 15900000 86239
ipcc 0 MSNBC 12050180260 6440000 77964
ipcc 0 BBC News 5099185254 15600000 20649
ipcc 0 Vox 3427290494 11900000 1677
ipcc 1 Fox News 16126544384 11000000 101812
ipcc 1 NBC News 6952022416 9650000 61006
ipcc 1 Sky News Australia 4140464784 3900000 137500
ipcc 1 PragerU 1898459165 3210000 3188
ipcc 2 Sky News Australia 4140464784 3900000 137500
ipcc 2 Sky News 4139448980 7370000 44104
ipcc 2 GBNews 1199821224 1170000 58097
ipcc 2 Rebel News 716906956 1640000 21112
ipcc 3 Fox News 16126544384 11000000 101812
ipcc 3 The Hill 1462188609 1900000 30762
ipcc 3 KTLA 5 1023133916 1020000 22849
ipcc 3 WUSA9 337193118 432000 52569
ipcc 4 WDR aktuell 83165245 103000 1139
ipcc 4 Carsten Jahn - TEAM HEIMAT 76120963 179000 879
ipcc 4 Stuttgarter Zeitung 60081195 60000 3424
ipcc 4 eingeSCHENKt.tv 19415430 131000 544
ipcc 5 De Nieuwe Wereld 26340396 101000 1797
ipcc 5 1000frolly PhD 10920557 30100 170
ipcc 5 Holland Gold 2880125 19800 181
ipcc 5 Tom Nelson 1587268 12400 219
refugees 0 Fox News 16126544384 11000000 101812
refugees 0 MSNBC 12050180260 6440000 77964
refugees 0 TEDx Talks 7712122864 40100000 210266
refugees 0 NBC News 6952022416 9650000 61006
refugees 1 CNN 15327838440 16000000 163405
refugees 1 ABC News 14083290454 15900000 86239
refugees 1 NBC News 6952022416 9650000 61006
refugees 1 BBC News 5099185254 15600000 20649
refugees 2 Fox News 16126544384 11000000 101812
refugees 2 MSNBC 12050180260 6440000 77964
refugees 2 NBC News 6952022416 9650000 61006
refugees 2 Sky News Australia 4140464784 3900000 137500
refugees 3 AllatRa TV English 2840110 29500 964
refugees 3 Creative Society 2664803 40600 1163
refugees 3 ALLATRA TV International 1507255 18500 982
refugees 4 VICE News 3156884335 8830000 6480
refugees 4 Global News 2490061489 4110000 37952
refugees 4 DW News 2241829454 5010000 33064
refugees 4 CBC News 1955874636 3450000 28005
refugees 5 TEDx Talks 7712122864 40100000 210266
refugees 5 NBC News 6952022416 9650000 61006
refugees 5 Democracy Now! 672210420 1980000 15630
Table S11: Top 4 most relevant YouTube channels by view count in the 5 largest communities.